### v1

Abstract
In this article, we present a modified variant of the Dai-Liao spectral conjugate gradient method, developed through an analysis of eigenvalues and inspired by a modified secant condition. We show that the proposed method is globally convergent for general nonlinear functions under standard assumptions. By incorporating the new secant condition and a quasi-Newton direction, we introduce updated spectral parameters. These changes ensure that the resulting search direction satisfies the sufficient descent property without relying on any line search. Numerical experiments show that the proposed algorithm performs better than several existing methods in terms of convergence speed and computational efficiency. Its effectiveness is further demonstrated through an application in signal processing.
keywords:
Spectral Conjugate Gradient Method Dai-Liao Method Optimization Signal Processing Secant Condition
MSC65K05 90C30 49M37
1 Introduction
[…]
find $\text{find }x^{*}\in\mathbb{R}^{m}\text{ such that }f(x^{*})=\min_{x\in\mathbb{R}^{m}}f(x).$ (1)
We denote the set of minimizers of $f$ as $\Omega_{\text{min}}(f,\mathbb{R}^{m})$.
The general iterative rule for solving (1) starts from an initial point $x_{0}\in\mathbb{R}^{m}$ and generates a sequence $\{x_{n}\}$ in $\mathbb{R}^{m}$ as follows:
$x_{n+1}=x_{n}+\alpha_{n}d_{n}\text{ for all }n\in\mathbb{N}_{0}=\mathbb{N}\cup\{0\},$ (2)
where the stepsize $\alpha_{n}$ is a positive real parameter determined via exact or inexact line search, $x_{n}$ is the previous iterative point, $x_{n+1}$ is the current point, and $d_{n}$ is a suitable search direction. The class of algorithms that follow the form (2) is known as line search algorithms. For such algorithms, only the search direction $d_{n}\in\mathbb{R}^{m}$ and the stepsize $\alpha_{n}\in(0,\infty)$ are required. An appropriate descent direction $d_{n}$ must be chosen such that it satisfies the descent condition:
$\langle\nabla f(x_{n}),d_{n}\rangle<0\text{ for all }n\in\mathbb{N}_{0}.$ (3)
A common choice for a descent direction is $d_{n}=-\nabla f(x_{n})$, which reduces the general line search iteration (2) to the gradient descent (GD) method:
$x_{n+1}=x_{n}-\alpha_{n}\nabla f(x_{n})\text{ for all }n\in\mathbb{N}_{0}.$
If there exists a constant $c>0$ such that
$\langle\nabla f(x_{n}),d_{n}\rangle\leq-c\|\nabla f(x_{n})\|^{2}\text{ for all}n\in\mathbb{N}_{0},$
then the vector $d_{n}$ is said to satisfy the sufficient descent condition.
Conjugate gradient (CG) methods are among the most effective algorithms for solving large-scale unconstrained optimization problems due to their low storage requirements. These methods require only first-order derivatives, making them highly suitable for large-scale applications. The first form of the CG method was introduced by Hestenes and Stiefel [1] in 1952 for solving linear systems of equations. Later, in 1964, Fletcher and Reeves [2] modified the original idea and developed the first nonlinear CG method for unconstrained optimization problems. Due to the strong theoretical foundation and excellent numerical performance of these algorithms, many variants of the CG method have been proposed (see, e.g., [3, 4, 5, 6]).
The successive approximations generated by the CG algorithms have the form (2), where $\alpha_{n}\in(0,\infty)$ is the step length determined through a line search along the CG direction $d_{n}$. The CG directions $d_{n}$ is defined as:
$d_{n}=\begin{cases}-\nabla f(x_{0}),&\text{ if }n=0,\\ -\nabla f(x_{n})+\beta_{n-1}d_{n-1},&\text{ if }n\geq 1,\end{cases}$ (4)
where $\beta_{n}$ is the CG parameter. There are many formulas for $\beta_{n}$, each resulting in different computational behavior [7].
Based on an extended conjugacy condition, one of the fundamental CG methods was proposed by Dai and Liao [4] with the CG parameter defined as:
$\beta_{n}^{DL}=\frac{\langle\nabla f(x_{n+1}),y_{n}\rangle}{\langle d_{n},y_{n}\rangle}-t\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},y_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0},$ (5)
where $t\in(0,\infty)$, $s_{n}=x_{n+1}-x_{n}$, and $y_{n}=\nabla f(x_{n+1})-\nabla f(x_{n})$. Note that when $t=0$, $\beta_{n}^{DL}$ simplifies to the CG parameter proposed by Hestenes and Stiefel [1]:
$\beta_{n}^{HS}=\frac{\langle\nabla f(x_{n+1}),y_{n}\rangle}{\langle d_{n},y_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0}.$
The CG parameter proposed by Hager and Zhang [8] is defined as:
$\beta_{n}^{HZ}=\frac{\langle\nabla f(x_{n+1}),y_{n}\rangle}{\langle d_{n},y_{n}\rangle}-2\frac{\|y_{n}\|^{2}}{\langle d_{n},y_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},y_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0},$
which can be viewed as an adaptive version of (5) for the specific choice $t=2\frac{\|y_{n}\|^{2}}{\langle d_{n},y_{n}\rangle}$. Similarly, the CG parameter proposed by Dai and Kou [9] is defined as:
$\beta_{n}^{DK}(\tau_{n})=\frac{\langle\nabla f(x_{n+1}),y_{n}\rangle}{\langle d_{n},y_{n}\rangle}-\left(\tau_{n}+\frac{\|y_{n}\|^{2}}{\langle s_{n},y_{n}\rangle}-\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}}\right)\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},y_{n}\rangle},$
where $\tau_{n}$ is a parameter corresponding to the scaling factor in the scaled memoryless Broyden-Fletcher-Goldfarb-Shanno (BFGS) method, which can be regarded as another adaptive form of (5).
Inspired by Shanno’s matrix perspective on CG algorithms [10], Babaie-Kafaki and Ghanbari [11] observed that the Dai-Liao search directions can be expressed as $d_{n+1}=-P_{n+1}\nabla f(x_{n+1})$ for all $n\in\mathbb{N}_{0}$, where $P_{n+1}=I-\frac{s_{n}y_{n}^{T}}{\langle s_{n},y_{n}\rangle}+t\frac{s_{n}s_{n}^{T}}{\langle s_{n},y_{n}\rangle}$ is a nonsingular matrix for $t\in(0,\infty)$ and $\langle s_{n},y_{n}\rangle\neq 0$; see [12]. Conducting an eigenvalue analysis on a symmetrized version of $P_{n+1}$, defined by $A_{n+1}=\frac{P_{n+1}+P_{n+1}^{T}}{2},$ we can derive a two-parameter formula for $t$ in the Dai-Liao CG parameter (5):
$t_{n}^{p,q}=p\frac{\|y_{n}\|^{2}}{\langle s_{n},y_{n}\rangle}-q\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}}\text{ for all }n\in\mathbb{N}_{0},$ (6)
which ensures the descent condition (3) for $p\in\left(\frac{1}{4},\infty\right)$ and $q\in\left(-\infty,\frac{1}{4}\right)$. However, in cases where $\frac{\|y_{n}\|^{2}}{\langle s_{n},y_{n}\rangle}$ and $\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}}$ are nearly equal, their subtraction may not be numerically stable [13]. Moreover, the parameter $t$ in CG parameter $\beta_{n}^{DL}$ must be nonnegative. To address these issues, they restricted $q$ in (6) to be nonpositive and proposed the following CG parameter:
$\beta_{n}^{DDL}=\frac{\langle\nabla f(x_{n+1}),y_{n}\rangle}{\langle d_{n},y_{n}\rangle}-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},y_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0},$
where $t_{n}^{p,q}$ is defined by (6) with $p\in\left(\frac{1}{4},\infty\right)$ and $q\in(-\infty,0]$. Note that for the parameter choice $(p,q)=(2,0)$ in (6), $\beta_{n}^{DDL}$ reduces to $\beta_{n}^{HZ}$. Similarly, if $(p,q)=(1,0)$, then $\beta_{n}^{DDL}$ is equivalent to $\beta_{n}^{DK}(\tau_{n})$ with the optimal choice of $\tau_{n}=\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}}$.
Another significant category of iterative methods used for solving unconstrained optimization problems is the spectral gradient method. Initially introduced by Barzilai and Borwein [14], these methods were later advanced by Raydan [15], who developed a spectral gradient approach specifically for large-scale unconstrained optimization challenges. The spectral conjugate direction is defined by
$d_{n}=\begin{cases}-\nabla f(x_{0}),&\text{ if }n=0,\\ -\theta_{n}\nabla f(x_{n})+\beta_{n-1}d_{n-1}&\text{ if }n\geq 1,\end{cases}$
where $\{\theta_{n}\}\subset(0,\infty)$ is spectral parameter and $\{\beta_{n}\}\subset(0,\infty)$ is conjugate parameter.
Building on the encouraging numerical results associated with spectral gradient methods [14, 15], Birgin and Martínez [16] proposed the integration of spectral gradient techniques with conjugate gradient methods, culminating in the development of the first spectral conjugate gradient (CG) method. The spectral CG methods represent a powerful enhancement of traditional CG methods by incorporating the principles of spectral gradients within the conjugate gradient framework. Furthering this line of research, Yu et al. [17] modified the spectral Perry’s CG method [18], resulting in an efficient spectral CG variant. Capitalizing on the robust convergence properties of quasi-Newton methods, Andrei [19] introduced a scaled BFGS preconditioned CG method in 2007, leveraging quasi-Newton techniques to enhance the effectiveness of conjugate gradient methods. This foundational idea has been elaborated upon by Andrei [20, 21, 22], as well as Babaie-Kafaki and Mahdavi-Amiri [23].
Using the modified secant condition proposed by Li and Fukushima [24], Zhou and Zhang [25] established the following modified secant condition:
$\nabla^{2}f(x_{n+1})s_{n}=z_{n},$
where
$z_{n}=y_{n}+h_{n}\|\nabla f(x_{n})\|^{r}s_{n},\quad h_{n}=\nu+\max\left\{-\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}},0\right\}\|\nabla f(x_{n})\|^{-r},$ (7)
with $r,\nu\in(0,\infty)$. The modified secant condition (7) plays a crucial role in establishing the global convergence of the MBFGS method for non-convex functions, see [25].
It is worth noting that the following relation holds independently of the line search:
$\langle z_{n},s_{n}\rangle$ $=\langle y_{n},s_{n}\rangle+\nu\|\nabla f(x_{n})\|^{r}\|s_{n}\|^{2}+\max\left\{-\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}},0\right\}\|s_{n}\|^{2}$
$\geq\nu\|\nabla f(x_{n})\|^{r}\|s_{n}\|^{2}>0\text{ for all }n\in\mathbb{N}_{0}.$
Inspired from the appropriate theoretical properties given by (7), Faramarzi and Amini [26] modified the spectral CG method proposed by Jian et al. [27] and considered the following conjugate parameter
$\beta_{n}^{ZDK}=\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-\frac{\|z_{n}\|^{2}}{\langle d_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0}.$
Based on this conjugate parameter, they construct the following spectral parameter
$\theta_{n+1}=\begin{cases}1-\frac{\|z_{n}\|^{2}}{\langle d_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle\nabla f(x_{n+1}),z_{n}\rangle},&\text{ if }1-\frac{\|z_{n}\|^{2}}{\langle d_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle\nabla f(x_{n+1}),z_{n}\rangle}\in\left[\frac{1}{4p}+|q|+\eta,\tau\right],\\ 1,&\text{ otherwise },\end{cases}$
where $\eta$ and $\tau$ are positive constants. To ensure the global convergence of this method, Faramarzi and Amini [26] used the line search satisfying the strong Wolfe conditions:
$f(x_{n}+\alpha_{n}d_{n})\leq f(x_{n})+\delta\alpha_{n}\langle\nabla f(x_{n}),d_{n}\rangle,$ (8)
$|\langle\nabla f(x_{n}+\alpha_{n}d_{n}),d_{n}\rangle|\leq-\sigma\langle\nabla f(x_{n}),d_{n}\rangle,$ (9)
where $10<\delta<\sigma<$ 1 and the following assumptions:
  1. (A1)
    The level set
    $C_{0}=\{x\in\mathbb{R}^{n}:f(x)\leq f(x_{0})\}$
    is bounded; namely, there exists a constant $M>0$ such that
    $\|x\|\leq M\text{ for all }x\in C_{0}.$
  2. (A2)
    In some neighbourhood $\Omega$ of $C_{0}$, $f$ is continuously differentiable and its gradient is Lipschitz continuous, namely there exists a constant $L>0$ such that
    $\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|\text{ for all }x,y\in\Omega.$ (10)
The primary motivation of this study is to introduce a modified spectral conjugate gradient method and to establish its global convergence for general functions under standard assumptions. This new method consistently satisfies the sufficient descent property, regardless of the line search employed. To demonstrate the effectiveness of the proposed approach, we present numerical results that validate its performance and show its superior convergence compared to existing methods. Furthermore, we explore its application in compressed sensing, emphasizing its practical relevance and significant advantages.
2 Preliminaries
In this section, we present some fundamental results from Euclidean space that will be used to establish the convergence of the proposed method.
Lemma 1.
([26]) Let $a,b,\sigma\in\mathbb{R}$. Then
$(c+\sigma d)^{2}\leq(1+\sigma^{2})(c^{2}+d^{2})\text{ for all }c,d,\sigma\geq 0,$
Lemma 2.
([26]) Let $\mathbb{R}^{m}$ be the Euclidean space. Then
$\langle u,v\rangle\leq\frac{1}{2}(\|u\|^{2}+\|v\|^{2})\text{ for all }u,v\in\mathbb{R}^{m}.$
Lemma 3.
Let $\mathbb{R}^{m}$ be the Euclidean space and let $\{u_{n}\}$ and $\{v_{n}\}$ be sequences in $\mathbb{R}^{m}$. Let $\{a_{n}\}$ and $\{b_{n}\}$ be sequences in $(0,\infty)$ and $u_{0}\in\mathbb{R}^{m}$ and $b_{0},\sigma\in(0,\infty)$ satisfying
  1. (C1)
    $a_{n}>a$ for all $n\in\mathbb{N}$,
  2. (C2)
    $\langle v_{n+1},u_{n+1}\rangle<0$ for all $n\in\mathbb{N}$,
  3. (C3)
    $|\langle v_{n+1},u_{n}\rangle|\leq-\sigma\langle v_{n},u_{n}\rangle$ for all $n\in\mathbb{N}$.
Suppose that
$u_{n+1}=-a_{n+1}v_{n+1}+b_{n}u_{n}\text{ for all }n\in\mathbb{N}.$ (11)
Then
$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}+a_{n+1}^{2}\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}\|v_{n+1}\|^{2}}\geq\frac{a}{1+\sigma^{2}}\frac{\|v_{n+1}\|^{2}}{\|u_{n}\|^{2}}.$
Proof.
From (11) and condition (C2), we obtain
$b_{n}^{2}\|u_{n}\|^{2}$ $=\|u_{n+1}\|^{2}+2a_{n+1}\langle v_{n+1},u_{n+1}\rangle+a_{n+1}^{2}\|v_{n+1}\|^{2}$
$\leq\|u_{n+1}\|^{2}+a_{n+1}^{2}\|v_{n+1}\|^{2},$
which implies that
$\frac{\|u_{n+1}\|^{2}}{\|u_{n}\|^{2}}\geq b_{n}^{2}-a^{2}\frac{\|v_{n+1}\|^{2}}{\|u_{n}\|^{2}}.$ (12)
Again, from (11), we have
$-\langle v_{n+1},u_{n+1}\rangle+b_{n}\langle v_{n+1},u_{n}\rangle=a_{n+1}\|v_{n+1}\|^{2}.$ (13)
From condition (C3), we have
$b_{n}\langle v_{n+1},u_{n}\rangle\leq b_{n}|\langle v_{n+1},u_{n}\rangle|\leq-\sigma b_{n}\langle v_{n},u_{n}\rangle\leq\sigma b_{n}|\langle v_{n},u_{n}\rangle|.$ (14)
From (13) and (14), we get
$-\langle v_{n+1},u_{n+1}\rangle+\sigma b_{n}|\langle u_{n},v_{n}\rangle|\geq a_{n+1}\|v_{n+1}\|^{2}.$
Squaring both side and applying Lemma 1 with $c=-\langle v_{n+1},u_{n+1}\rangle>0$ (from (C2)) and $d=b_{n}|\langle v_{n},u_{n}\rangle|>0$ gives
$\langle v_{n+1},u_{n+1}\rangle^{2}+b_{n}^{2}\langle v_{n},u_{n}\rangle^{2}\geq\frac{a_{n+1}^{2}}{1+\sigma^{2}}\|v_{n+1}\|^{4}.$
Since $a_{n+1}>a$, this yields that
$\langle v_{n+1},u_{n+1}\rangle^{2}+b_{n}^{2}\langle v_{n},u_{n}\rangle^{2}$ $\geq\frac{a}{1+\sigma^{2}}\|v_{n+1}\|^{4},$
which implies that
$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}$ $=\frac{1}{\|u_{n+1}\|^{2}}\left[\langle v_{n+1},u_{n+1}\rangle^{2}+\frac{\|u_{n+1}\|^{2}}{\|u_{n}\|^{2}}\langle v_{n},u_{n}\rangle^{2}\right]$
$\geq\frac{1}{\|u_{n+1}\|^{2}}\left[\frac{a}{1+\sigma^{2}}\|v_{n+1}\|^{4}+\left(\frac{\|u_{n+1}\|^{2}}{\|u_{n}\|^{2}}-b_{n}^{2}\right)\langle v_{n},u_{n}\rangle^{2}\right].$ (15)
From (12) and (15), we conclude that
$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n+1},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}\geq\frac{\|v_{n+1}\|^{4}}{\|u_{n+1}\|^{2}}\left[\frac{a}{1+\sigma^{2}}-a_{n+1}^{2}\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}\|v_{n+1}\|^{2}}\right],$
which implies that
$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}+a_{n+1}^{2}\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}\|v_{n+1}\|^{2}}\geq\frac{a}{1+\sigma^{2}}\frac{\|v_{n+1}\|^{2}}{\|u_{n}\|^{2}}.$
Lemma 4.
([28]) Let $f:\mathbb{R}^{n}\to\mathbb{R}$ such that assumptions (A1) and (A2) hold. Consider any iterative method of the form (2), where $d_{n}$ is a descent direction and $\alpha_{n}$ is obtained by the strong Wolfe conditions (8) and (9). Then
$\sum_{n=0}^{\infty}\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}<\infty.$
Remark 5.
From Lemma 4, we have
$\lim_{n\to\infty}\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}=0.$ (16)
3 A modified descent Dai-Liao spectral conjugate gradient method and its convergence analysis
In this section, inspired by the concepts from quasi-Newton methods, we introduce an effective modified spectral conjugate gradient (CG) method.
Motivated by the theoretical properties provided by (7) and the Dai-Liao conjugate gradient parameter (5), we propose the following modified Dai-Liao conjugate parameter:
$\beta_{n}=\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-t\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0},$ (17)
where $t$ is a nonnegative parameter.
Note that, by (4) and (17), the search direction can be expressed as
$d_{n+1}=-Q_{n+1}\nabla f(x_{n+1})\text{ for all }n\in\mathbb{N}_{0},$
where
$Q_{n+1}=I-\frac{s_{n}z_{n}^{T}}{\langle s_{n},z_{n}\rangle}+t\frac{s_{n}s_{n}^{T}}{\langle s_{n},y_{n}\rangle}.$
Following the methodology in [11], we propose the following formula for the parameter $t$ in the modified Dai-Liao conjugate parameter (17):
$t_{n}^{p,q}=p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\text{ for all }n\in\mathbb{N}_{0},$ (18)
with $p\in\left(\frac{1}{4},\infty\right)$ and $q\in\left(-\infty,\frac{1}{4}\right)$. Using this parameter $t_{n}^{p,q}$, we define the following modified descent Dai-Liao (MDDL) conjugate parameter:
$\beta_{n}^{MDDL}=\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0}.$ (19)
Now, we consider the following spectral conjugate gradient search direction:
$d_{n}=\begin{cases}-\nabla f(x_{0}),&\text{ if }n=0,\\ -\theta_{n}\nabla f(x_{n})+\beta_{n-1}^{MDDL}d_{n-1}&\text{ if }n\geq 1,\end{cases}$ (20)
where $\{\theta_{n}\}$ is a sequence in $(0,\infty)$.
Pproposition 6.
Let $f:\mathbb{R}^{m}\to\mathbb{R}$ be a differentiable function and let $\{d_{n}\}$ be a sequence of spectral CG search directions generated by (20). Then we have the following:
  • (a)
    For all $n\in\mathbb{N}_{0},$ we have
    $\langle\nabla f(x_{n}),d_{n}\rangle\leq-\kappa_{n}\|\nabla f(x_{n})\|^{2}$ (21)
    where
    $\kappa_{n}=\left\{\begin{array}[]{ll}1&\text{if }n=0,\\ \theta_{n}-\left(\frac{1}{4p}+|q|\right)&\text{if }n\geq 1.\end{array}\right.$
  • (b)
    If $\theta_{n}>\frac{1}{4p}+|q|$ for all $n\in\mathbb{N}$, then the direction given by (20) is a descent direction.
Proof.
(a) Note that $\langle\nabla f(x_{0}),d_{0}\rangle=-\|\nabla f(x_{0})\|^{2}$. Then the inequality (21) holds for $n=0$. We show that the inequality (21) holds for $n>0$. Let $n\in\mathbb{N}_{0}$. Note that $p\in\left(\frac{1}{4},\infty\right)$ and $q\in\left(-\infty,\frac{1}{4}\right)$. From (18), (19) and (20), we get
$\langle\nabla f(x_{n+1}),d_{n+1}\rangle$
$=\langle\nabla f(x_{n+1}),-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}^{MDDL}d_{n}\rangle$
$=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad-\frac{t_{n}^{p,q}\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}$
$=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad-\left(p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\right)\frac{\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}$
$=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
$\quad-p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}\frac{\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}+q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\frac{\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}.$
Since $s_{n}=\alpha_{n}d_{n}$, we have
$\langle\nabla f(x_{n+1}),d_{n+1}$
$=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}-p\frac{\|z_{n}\|^{2}\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle}$
$\quad+q\frac{\langle d_{n},z_{n}\rangle}{\|d_{n}\|^{2}}\frac{\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle}$
$=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}-p\frac{\|z_{n}\|^{2}\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle^{2}}$
$\quad+q\frac{\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\|d_{n}\|^{2}}.$ (22)
From Lemma 2 with $u_{n}=\frac{1}{\sqrt{2p}}\nabla f(x_{n+1})$ and $v_{n}=\frac{\sqrt{2p}\langle\nabla f(x_{n+1}),d_{n}\rangle z_{n}}{\langle d_{n},z_{n}\rangle},$ we obtain
$\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}\leq\frac{1}{4p}\|\nabla f(x_{n+1})\|^{2}+\frac{p\|z_{n}\|^{2}\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle^{2}}.$ (23)
On the other hand, by Cauchy-Schwarz inequality, we have
$q\frac{\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\|d_{n}\|^{2}}\leq|q|\frac{\|d_{n}\|^{2}\|\nabla f(x_{n+1})\|^{2}}{\|d_{n}\|^{2}}=|q|\|\nabla f(x_{n+1})\|^{2}.$ (24)
Combining (22), (23) and (24), we conclude that
$\langle\nabla f(x_{n+1}),d_{n+1}\rangle$ $\leq-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{1}{4p}\|\nabla f(x_{n+1})\|^{2}+|q|\|\nabla f(x_{n+1})\|^{2}$
$=-\left(\theta_{n+1}-\left(\frac{1}{4p}+|q|\right)\right)\|\nabla f(x_{n+1})\|^{2}.$
(b) Suppose that $\theta_{n}>\frac{1}{4p}+|q|$ for all $n\in\mathbb{N}$. Then, from (21), we get $\langle\nabla f(x_{n}),d_{n}\rangle<0$ for all $n\in\mathbb{N}_{0}$. Therefore, $\{d_{n}\}$ is a descent direction. ∎
Now, we construct the spectral parameter for our conjugate gradient method based on the conjugate parameter $\beta_{n}^{MDDL}$.
Quasi-Newton methods are a prominent class of iterative optimization techniques that do not require computing the exact Hessian matrix at each iteration. The primary motivation behind these methods is to achieve the rapid convergence rate of the classical Newton method while reducing the computational cost. Instead of calculating the Hessian matrix directly, quasi-Newton methods construct a sequence of Hessian approximations. Inspired by the quasi-Newton direction, the search direction $d_{n+1}$ defined by (20) is designed to mimic the behavior of the quasi-Newton method. Consequently, we aim to determine a suitable spectral parameter $\theta_{n+1}$ such that
$-B_{n+1}^{-1}\nabla f(x_{n+1})\approx-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}^{MDDL}d_{n},$
where $B_{n+1}$ is an approximation of the Hessian matrix $\nabla^{2}f(x_{n+1})$. Taking inner product with $B_{n+1}s_{n}$ gives that
$-\langle s_{n},\nabla f(x_{n+1})\rangle\approx-\theta_{n+1}\langle B_{n+1}s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle B_{n+1}s_{n},d_{n}\rangle,$
which implies that
$\theta_{n+1}\approx\frac{1}{\langle B_{n+1}s_{n},\nabla f(x_{n+1})\rangle}(\langle s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle B_{n+1}s_{n},d_{n}\rangle).$
Since $B_{n+1}$ must satisfies the modified secant condition $B_{n+1}s_{n}=z_{n}$, we have
${\theta_{n+1}}$
$=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}(\langle s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle z_{n},d_{n}\rangle)$
$=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}\left(\langle s_{n},\nabla f(x_{n+1})\rangle+\left(\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right)\langle z_{n},d_{n}\rangle\right)$
$=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}\left(\langle s_{n},\nabla f(x_{n+1})\rangle+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle z_{n},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle\langle z_{n},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right)$
$=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}\left(\langle s_{n},\nabla f(x_{n+1})\rangle+\langle\nabla f(x_{n+1}),z_{n}\rangle-t_{n}^{p,q}\langle\nabla f(x_{n+1}),s_{n}\rangle\right)$
$=1-\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}\left({t_{n}^{p,q}\langle s_{n},\nabla f(x_{n+1})\rangle}-\langle s_{n},\nabla f(x_{n+1})\rangle\right)$ (25)
$=1-(t_{n}^{p,q}-1)\frac{\langle s_{n},\nabla f(x_{n+1})\rangle}{\langle z_{n},\nabla f(x_{n+1})\rangle}.$ (26)
We neglect $\langle s_{n},\nabla f(x_{n+1})\rangle$ in (25), which leads to
${\theta_{n+1}}=1-\frac{t_{n}^{p,q}\langle s_{n},\nabla f(x_{n+1})\rangle}{\langle z_{n},\nabla f(x_{n+1})\rangle}.$ (27)
We denote the $\theta_{n+1}$ obtained in equation (26) by $\theta_{n+1}^{N+}$ and the $\theta_{n+1}$ obtained in equation (27) by $\theta_{n+1}^{N-}$. To obtain the sufficient descent condition and the bounded property of the spectral parameters, we propose the following spectral parameters:
$\theta_{n+1}=\begin{cases}\theta_{n+1}^{R},&\text{ if }\theta_{n+1}^{R}\in\left[\frac{1}{4p}+|q|+\eta,\tau\right],\\ 1,&\text{ otherwise }\end{cases}$ (28)
or
$\theta_{n+1}=\begin{cases}\theta_{n+1}^{N},&\text{ if }\theta_{n+1}^{N}\in\left[\frac{1}{4p}+|q|+\eta,\tau\right],\\ 1,&\text{ otherwise },\end{cases}$ (29)
where $\eta$ and $\tau$ are positive constants. Relation (28) and (29) lead to
$\frac{1}{4p}+|q|+\eta\leq\theta_{n+1}\leq\tau\textrm{ for all }n\in\mathbb{N}_{0}.$ (30)
Now, we propose the following modified descent Dai-Liao spectral conjugate gradient method (MDDLSCG):
Input: Let $x_{0}\in\mathbb{R}^{n}$ and let $\epsilon,\delta,\sigma,\eta,\tau,r,\nu\in(0,\infty)$ such that $10<\delta<\sigma<$ 1.
Step 0: Set $d_{0}=-\nabla f(x_{0})$.
Step 1: If $\|\nabla f(x_{n})\|_{\infty}<\epsilon$, stop.
Step 2: Determine the step length $\alpha_{n}$ satisfying the strong Wolfe conditions (8) and (9).
Step 3: Set $x_{n+1}=x_{n}+\alpha_{n}d_{n}$.
Step 4: Compute $d_{n+1}$ by
$d_{n+1}=-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}d_{n}\text{ for all }n\in\mathbb{N}_{0},$
where CG parameter $\beta_{n}^{MDDL}$ is given by (19) and spectral parameter is given by (28) or (29).
Step 5: Set $n=n+1$ and return to Step 1.
Algorithm 1 MDDLSCG Algorithm
Remark 7.
(i) Using $\theta_{n}\geq\frac{1}{4p}+|q|+\eta$ and (21), we conclude that
$\langle\nabla f(x_{n}),d_{n}\rangle\leq-\eta\|\nabla f(x_{n})\|^{2}\text{ for all }n\in\mathbb{N}.$ (31)
(ii) If $\|\nabla f(x_{n})\|\neq 0$, then, from (9) and (31), we see that
$\langle d_{n},y_{n}\rangle\geq-(1-\sigma)\langle\nabla f(x_{n}),d_{n}\rangle\geq(1-\sigma)\eta\|\nabla f(x_{n})\|^{2}>0,$ (32)
which implies, from (7) that $h_{n}=\nu$ and hence
$z_{n}=y_{n}+\nu\|\nabla f(x_{n})\|^{r}s_{n},$ (33)
where $r,\nu\in(0,\infty)$.
Now, we analyze the global convergence of the MDDLSCG Algorithm.
Remark 8.
  1. (i)
    From (8) and (31), we can see that $\{f(x_{n})\}$ is decreasing. Thus, the sequence $\{x_{n}\}$ generated by Algorithm 1 is contained in $C_{0}$.
  2. (ii)
    From Assumption (A2), we have
    $\|\nabla f(x)\|$ $\leq\|\nabla f(x)-\nabla f(x_{0})\|+\|\nabla f(x_{0})\|$
    $\leq L\|x-x_{0}\|+\|\nabla f(x_{0})\|$
    $\leq 2ML+\|\nabla f(x_{0})\|:=\gamma\text{ for all }x\in C_{0}.$ (34)
In addition, suppose that $\nabla f(x_{n})\neq 0$ for all $n\in\mathbb{N}_{0}$; otherwise, a stationary point will be found.
Lemma 9.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a function and let $\{x_{n}\}$ be sequence generated by Algorithm 1 such that Assumptions (A1) and (A2) hold. Suppose that $\theta_{n}>\frac{1}{4p}+|q|$ for all $n\in\mathbb{N}$. If $\liminf_{n\to\infty}\|\nabla f(x_{n})\|\neq 0$, then $\sum_{n=0}^{\infty}\frac{\|\nabla f(x_{n})\|^{4}}{\|d_{n}\|^{2}}<\infty.$
Proof.
Since $\liminf_{n\to\infty}\|\nabla f(x_{n})\|\neq 0$. Then, there exists $\epsilon>0$ such that $\|\nabla f(x_{n})\|\geq\epsilon$ for all $n\in\mathbb{N}_{0}$. This together with (16) implies that
$\lim_{n\to\infty}\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}\|\nabla f(x_{n+1})\|^{2}}=0.$ (35)
Let $u_{n}=d_{n}$, $v_{n}=\nabla f(x_{n})$, $a_{n}=\theta_{n}$ and $b_{n}=\beta_{n}$ in Lemma 3. Since $\theta_{n}>\frac{1}{4p}+|q|$, it follows from Remark 7 that $\langle\nabla f(x_{n}),d_{n}\rangle<0$ for all $n\in\mathbb{N}_{0}$. Also from strong Wolfe condition (9), we have $|\langle\nabla f(x_{n+1}),d_{n}\rangle|\leq-\sigma\langle\nabla f(x_{n}),d_{n}\rangle$. Then, clearly all the conditions (C1), (C2) and (C3) of Lemma 3 is satisfied. Thus, we have
$\frac{\langle\nabla f(x_{n+1}),d_{n+1}\rangle^{2}}{\|d_{n+1}\|^{2}}+\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}\geq b_{1}\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}-\theta_{n+1}^{2}\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}\|\nabla f(x_{n+1})\|^{2}},$
where $b_{1}=\left(\frac{1}{4p}+|q|\right)\frac{1}{1+\sigma^{2}}$. It follows from (35) that for all sufficiently large $n$, we have
$\frac{\langle\nabla f(x_{n+1}),d_{n+1}\rangle^{2}}{\|d_{n+1}\|^{2}}+\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}\geq b_{1}\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}.$ (36)
From (36) and Lemma 4, we get $\sum_{n=1}^{\infty}\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}<\infty.$
Theorem 10.
Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a function and let $\{x_{n}\}$ be sequence generated by Algorithm 1 such that Assumptions (A1) and (A2) hold. Then $\liminf_{n\to\infty}\|\nabla f(x_{n})\|=0.$
Proof.
Suppose the claim of the theorem is not true, i.e., there exists a positive constant $\xi$ such that
$\|\nabla f(x_{n})\|>\xi\text{ for all }n\in\mathbb{N}_{0}.$ (37)
From (32), (33) and (37), we obtain
$\langle d_{n},z_{n}\rangle=\langle d_{n},y_{n}\rangle+\nu\|\nabla f(x_{n})\|^{r}\alpha_{n}\|d_{n}\|^{2}\geq\nu\|\nabla f(x_{n})\|^{r}\alpha_{n}\|d_{n}\|^{2}\geq\nu\xi^{r}\alpha_{n}\|d_{n}\|^{2}.$ (38)
From (33), (10) and ((ii)), we have
$\|z_{n}\|\leq\|y_{n}\|+\nu\|\nabla f(x_{n})\|^{r}\|s_{n}\|\leq L\|s_{n}\|+\nu\gamma^{r}\|s_{n}\|=(L+\nu\gamma^{r})\|s_{n}\|.$ (39)
From the Lipschitz continuity of the $\nabla f$ and (39), we obtain
$t_{n}^{p,q}\leq\left|p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\right|\leq p\frac{(L+\nu\gamma^{r})^{2}\|s_{n}\|^{2}}{\mu\|s_{n}\|^{2}}+|q|\frac{(L+\nu\gamma^{r})\|s_{n}\|^{2}}{\|s_{n}\|^{2}}=p\frac{L^{2}}{\mu}+L|q|.$
Thus, the relations (19), (10), ((ii)), (38), (39), along with Cauchy-Schwarz inequality, conclude
$\|\beta_{n}^{MDDL}\|$ $=\left\|\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
$=\left\|\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-\left(p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\right)\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
$=\left\|\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}+q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
$=\left\|\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|+p\left\|\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|+|q|\left\|\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
$=\frac{\|\nabla f(x_{n+1})\|\|z_{n}\|}{\langle d_{n},z_{n}\rangle}+p\frac{\|z_{n}\|^{2}\|\nabla f(x_{n+1})\|\|d_{n}\|}{\langle d_{n},z_{n}\rangle^{2}}+|q|\frac{\|\nabla f(x_{n+1})\|}{\|d_{n}\|}$ (40)
$\leq\frac{\gamma(L+\nu\gamma^{r})}{\nu\xi^{r}\|d_{n}\|}+\frac{p\gamma(L+\nu\gamma^{r})^{2}}{\nu^{2}\xi^{2r}\|d_{n}\|}+|q|\frac{\gamma}{\|d_{n}\|}.$
Note that $\theta_{n+1}\leq\tau$ by (30). Hence, from (20), (40), we get
$\|d_{n+1}\|$ $=\|-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}^{MDDL}d_{n}\|$
$\leq\|\theta_{n+1}\|\|\nabla f(x_{n+1})\|+\|\beta_{n}^{MDDL}\|\|d_{n}\|$
$\leq\tau\gamma+\frac{\gamma(L+\nu\gamma^{r})}{\nu\xi^{r}}+\frac{p\gamma(L+\nu\gamma^{r})^{2}}{\nu^{2}\xi^{2r}}+|q|\gamma.$ (41)
From (41), we observe that
$\frac{1}{\|d_{n+1}\|^{2}}\geq\frac{1}{E^{2}}\textrm{ for all }n\in\mathbb{N}_{0},$
where $E=\tau\gamma+\frac{\gamma(L+\nu\gamma^{r})}{\nu\xi^{r}}+\frac{p\gamma(L+\nu\gamma^{r})^{2}}{\nu^{2}\xi^{2r}}+|q|\gamma$. This implies that
$\sum_{n=0}^{\infty}\frac{1}{\|d_{n}\|^{2}}=\infty.$ (42)
From (37) and Lemma 9, we have
$\sum_{n=0}^{\infty}\frac{1}{\|d_{n}\|^{2}}=\sum_{n=0}^{\infty}\frac{\|\nabla f(x_{n})\|^{4}}{\|d_{n}\|^{2}\|\nabla f(x_{n})\|^{4}}\leq\frac{1}{\xi^{4}}\sum_{n=0}^{\infty}\frac{\|\nabla f(x_{n})\|^{4}}{\|d_{n}\|^{2}}<\infty,$
which contradicts (42). Therefore, $\liminf_{n\to\infty}\|\nabla f(x_{n})\|=0$. ∎
4 Numerical Experiments
In this section, we discuss the numerical performance of our proposed method (MDDLSCG), comparing it with Algorithm 3.1 in [26] (denoted by MSCG) and Scaled conjugate gradient method (denoted by ScCGM) in [29]. All numerical experiments were conducted in MATLAB R2024b on an HP workstation with Intel Core i7-13700 processor, 32 GB RAM and Window 11.
During the numerical experiments, we set the parameters $\delta=0.01$, $\sigma=0.1$ for all the algorithms, $\eta=0.001$, $\tau=10$, $r=1$ and $\nu=0.001$ for algorithms MSCG and MDDLSCG and $p=0.4$ and $q=0.2$ for the algorithm MDDLSCG. During the implementation, the iterations are terminated, when it meets the stopping criterion $E_{n}=\|\nabla f(x_{n})\|_{\infty}<10^{-15}$.
For the numerical experiment, we consider the test problem of finding the minima of Beale function. The Beale function is multimodal with sharp peaks at the corners of the input domain, which is defined from the general Beale function as follows:
$f(x,y)=\left(1.5-x+xy\right)^{2}+\left(2.25-x+xy^{2}\right)^{2}+\left(2.625-x+xy^{3}\right)^{2}\text{ for all }(x,y)\in\mathbb{R}^{2}.$
The global minimum is located at $(x,y)=(3,0.5)$, where $f(x,y)=0$.
The numerical results of all the algorithms (MDDLSCG, MSCG and ScCG) are summarized in Table 1, while their performance is illustrated in Figure LABEL:bealee. In Figure LABEL:bealee (a), the iteration paths are visualized on the contour plot of the Beale function, showing how each method converges to the minimizer. Figure LABEL:bealee (b) depicts the convergence rate by plotting the norm of the gradient, $\|\nabla f(x_{n})\|_{\infty}$ versus the number of iterations on a logarithmic scale.
From the data presented, we can clearly observe that our method, MDDLSCG, outperforms the MSCG algorithm and ScCG in all tested scenarios. This superiority is evidenced by the metrics shown in the table and figures, which highlight the efficiency and effectiveness of MDDLSCG compared to MSCG and ScCG.
Method itr Tcpu $E_{n}$
MDDLSCG 21 5.957900e-03 3.580469e-15
MSCG 30 1.783450e-02 4.662937e-15
ScCG 139 9.749100e-03 8.840151e-15
Table 1: Comparison table for MDDLSCG, MSCG and ScCG for Beale problems
5 Application in compressed sensing
Compressed sensing is a well-known signal processing strategy for efficiently acquiring and reconstructing signals with a sparse structure, allowing for memoryless and compact storage of the signals [30]. The importance of compressed sensing in real-world applications, such as machine learning, compressive imaging, radar, wireless sensor networks, medical imaging, astrophysical signals, and video coding, has been highlighted in [31].
The compressed sensing problem involves finding a sparse solution to a highly underdetermined system, represented as $Ax=b$, where $A\in\mathbb{R}^{m\times n}$ (with $m\ll n$) and $b\in\mathbb{R}^{m}$ [30]. A common approach is to solve the following optimization problem:
$f(x)=\frac{1}{2}\|Ax-b\|^{2}+\mu\phi(x),$
where $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is a penalty function, and $\mu>0$ is a parameter that controls the trade-off between sparsity and reconstruction fidelity. A popular choice for $\phi$ is $\phi(x)=\|x\|_{1}$, leading to the well-known Basis Pursuit Denoising (BPD) problem, which has been extensively studied in the literature; see [32] and the references therein. However, solving BPD is challenging due to the non-smoothness of the $\ell_{1}$-penalty term. To address this, Zhu et al. [33] proposed a relaxation of BPD using Nesterov’s smoothing technique [34], resulting in the following formulation:
$f_{\lambda}(x)=\sum_{i=1}^{n}\psi_{\lambda}(|x_{i}|),$
where $x_{i}$ represents the $i$-th component of the vector $x$, and $\psi_{\lambda}$ is defined as:
$\psi_{\lambda}(|x_{i}|)=\max_{\gamma\in[0,1]}\left\{\gamma|x_{i}|-\frac{1}{2}\lambda\gamma^{2}\right\}$
which simplifies to:
$\psi_{\lambda}(|x_{i}|)=\begin{cases}\frac{|x_{i}|^{2}}{2\lambda},&\text{if }|x_{i}|<\lambda,\\ |x_{i}|-\frac{\lambda}{2},&\text{if }|x_{i}|\geq\lambda.\end{cases}$
Here, $f_{\lambda}$ is known as the Huber function [35], and its gradient $\nabla f_{\lambda}$ has a Lipschitz constant of $\frac{1}{\lambda}$ [36], with the explicit form given by:
$(\nabla f_{\lambda})_{i}=\begin{cases}\frac{x_{i}}{\lambda},&\text{if }|x_{i}|<\lambda,\\ \text{sgn}(x_{i}),&\text{otherwise},\end{cases}$
where $\lambda>0$ is a constant and $\text{sgn}(\cdot)$ denotes the signum function.
Therefore, instead of solving the original problem $f(x)=\frac{1}{2}\|Ax-b\|^{2}+\mu\|x\|_{1}$, we solve the smoothed version:
$\min_{x\in\mathbb{R}^{n}}f(x)=\frac{1}{2}\|Ax-b\|^{2}+\mu f_{\lambda}(x).$
Clearly, the problem above is an unconstrained smooth convex optimization problem, and its gradient is given by:
$\nabla f(x)=\lambda\nabla f_{\lambda}(x)+A^{T}(Ax-b).$
For the numerical experiments, we aim to reconstruct a sparse signal of length $n$ from $m$-length observations ($m\ll n$). The parameters for the MDDLSCG, MSCG and ScCG algorithms are set as follows: $\mu=\max\{2^{-7},0.001\|A^{T}b\|_{\infty}\}$, $\lambda=\min\{0.001,0.048\|A^{T}b\|_{\infty}\}$, $\sigma=0.1$, $\delta=0.01$, $\eta=0.001$, $\tau=10$, $r=1$ and $\nu=0.001$.
[…]
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(x_{i}-x_{i}^{*})^{2}\leq 10^{-5}.$
To evaluate the restoration performance qualitatively, we report the relative error (RelErr) of the recovered signal, defined as:
$\text{RelErr}=\frac{\|x^{*}-x\|}{\|x\|}.$
Let $k$ denote the number of nonzero elements in the original signal. For a fair comparison, we consider two different settings: $(m,n,k)=(128,512,16)$, $(256,1024,32)$.
The detailed numerical results are presented in Table 2, where we report the average number of iterations (Itr), average CPU time (Tcpu), mean squared error (MSE), and relative error (RelErr) for each setting. The graph depicting the relationship between relative error and the number of iterations is shown in Figures LABEL:stpp512(a) and LABEL:stopp(a), while the graph illustrating MSE versus the number of iterations is presented in Figures LABEL:stpp512(b) and LABEL:stopp(b). Additionally, the recovery of the signal is visualized in Figures 4 and 6.
From Table 2, Figures LABEL:stpp512, and LABEL:stopp, it is clear that the MDDLSCG algorithm surpasses the MSCG and ScCG algorithms in performance across all cases.
Figure 4: Signal recovery for data $(m,n,k)=(128,512,16)$
Figure 6: Signal recovery for data $(m,n,k)=(256,1024,32)$
Dimensions $(m,n,k)$ Method Itr Tcpu MSE RelErr
(128,512,16) MDDLSCG 272 1.186382e-01 2.033486e-06 7.487872e-03
MSCG 493 1.606987e-01 2.033495e-06 7.488087e-03
ScCG 3653 1.135658e+00 2.033588e-06 7.712389e-03
(256,1024,32) MDDLSCG 291 2.163503e-01 2.128598e-06 9.721181e-03
MSCG 1405 8.339790e-01 2.128516e-06 9.720993e-03
ScCG 3002 1.746555e+00 2.229608e-04 9.949161e-02
Table 2: Comparison table for methods MDDLSCG, MSCG and ScCG
6 Conclusions
We presented a modified descent Dai-Liao spectral conjugate gradient (MDDLSCG) method that incorporates a new secant condition and quasi-Newton direction. The updated method introduces spectral parameters ensuring a sufficient descent property independent of line search techniques. Theoretical analysis establishes the global convergence of the method for general nonlinear functions under standard assumptions. Numerical experiments and an application in signal processing demonstrate the improved performance and practical effectiveness of the proposed algorithm, highlighting its potential in optimization and related applications.

### v2

@@
-
+Modified Dai-Liao Spectral Conjugate Gradient Method with Application to Signal Processing
 Abstract
@@
 keywords:
-Spectral Conjugate Gradient Method Dai-Liao Method Optimization Signal Processing Secant Condition
-MSC65K05 90C30 49M37
+Spectral Conjugate Gradient Method , Dai-Liao Method , Optimization , Signal Processing
+MSC:
+[2020] 65K05 , 90C30 , 49M37
+[…]
+organization=Department of Mathematics, Banaras Hindu University,city=Varanasi, country=India
+[…]
+organization=Department of Mathematics, The Technion- Israel Institute of Technology,city=Haifa, country=Israel
+[…]
+organization=Department of Applied Mathematics and Scientific Computing, Indian Institute of Technology Roorkee,country=India
 1 Introduction
@@
 $x_{n+1}=x_{n}+\alpha_{n}d_{n}\text{ for all }n\in\mathbb{N}_{0}=\mathbb{N}\cup\{0\},$ (2)
-where the stepsize $\alpha_{n}$ is a positive real parameter determined via exact or inexact line search, $x_{n}$ is the previous iterative point, $x_{n+1}$ is the current point, and $d_{n}$ is a suitable search direction.
+where the stepsize $\alpha_{n}$ is a positive real parameter determined via exact or inexact line search, $x_{n}$ is the previous iterative point, $x_{n+1}$ is the current point and $d_{n}$ is a suitable search direction.
 The class of algorithms that follow the form (2) is known as line search algorithms.
@@
 Inspired by Shanno’s matrix perspective on CG algorithms [10], Babaie-Kafaki and Ghanbari [11] observed that the Dai-Liao search directions can be expressed as $d_{n+1}=-P_{n+1}\nabla f(x_{n+1})$ for all $n\in\mathbb{N}_{0}$, where $P_{n+1}=I-\frac{s_{n}y_{n}^{T}}{\langle s_{n},y_{n}\rangle}+t\frac{s_{n}s_{n}^{T}}{\langle s_{n},y_{n}\rangle}$ is a nonsingular matrix for $t\in(0,\infty)$ and $\langle s_{n},y_{n}\rangle\neq 0$; see [12].
-Conducting an eigenvalue analysis on a symmetrized version of $P_{n+1}$, defined by $A_{n+1}=\frac{P_{n+1}+P_{n+1}^{T}}{2},$ we can derive a two-parameter formula for $t$ in the Dai-Liao CG parameter (5):
+Conducting an eigenvalue analysis on a symmetrized version of $P_{n+1}$, defined by $A_{n+1}=\frac{P_{n+1}+P_{n+1}^{T}}{2}$, we can derive a two-parameter formula for $t$ in the Dai-Liao CG parameter (5), given by
 $t_{n}^{p,q}=p\frac{\|y_{n}\|^{2}}{\langle s_{n},y_{n}\rangle}-q\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}}\text{ for all }n\in\mathbb{N}_{0},$ (6)
@@
 The spectral conjugate direction is defined by
-$d_{n}=\begin{cases}-\nabla f(x_{0}),&\text{ if }n=0,\\ -\theta_{n}\nabla f(x_{n})+\beta_{n-1}d_{n-1}&\text{ if }n\geq 1,\end{cases}$
+$d_{n}=\begin{cases}-\nabla f(x_{0}),&\text{ if }n=0,\\ -\theta_{n}\nabla f(x_{n})+\beta_{n-1}d_{n-1},&\text{ if }n\geq 1,\end{cases}$
 where $\{\theta_{n}\}\subset(0,\infty)$ is spectral parameter and $\{\beta_{n}\}\subset(0,\infty)$ is conjugate parameter.
@@
 where
-$z_{n}=y_{n}+h_{n}\|\nabla f(x_{n})\|^{r}s_{n},\quad h_{n}=\nu+\max\left\{-\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}},0\right\}\|\nabla f(x_{n})\|^{-r},$ (7)
-with $r,\nu\in(0,\infty)$. The modified secant condition (7) plays a crucial role in establishing the global convergence of the MBFGS method for non-convex functions, see [25].
+$z_{n}=y_{n}+h_{n}\|\nabla f(x_{n})\|^{r}s_{n},\quad h_{n}=\nu+\max\left\{-\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}},0\right\}\|\nabla f(x_{n})\|^{-r}$ (7)
+with $r,\nu\in(0,\infty)$. The modified secant condition (7) plays a crucial role in establishing the global convergence of the MBFGS method for non-convex functions; see [25].
 It is worth noting that the following relation holds independently of the line search:
@@
 $\beta_{n}^{ZDK}=\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-\frac{\|z_{n}\|^{2}}{\langle d_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0}.$
-Based on this conjugate parameter, they construct the following spectral parameter
+Based on this conjugate parameter, they construct the following spectral parameter:
 $\theta_{n+1}=\begin{cases}1-\frac{\|z_{n}\|^{2}}{\langle d_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle\nabla f(x_{n+1}),z_{n}\rangle},&\text{ if }1-\frac{\|z_{n}\|^{2}}{\langle d_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle\nabla f(x_{n+1}),z_{n}\rangle}\in\left[\frac{1}{4p}+|q|+\eta,\tau\right],\\ 1,&\text{ otherwise },\end{cases}$
@@
 The primary motivation of this study is to introduce a modified spectral conjugate gradient method and to establish its global convergence for general functions under standard assumptions.
-This new method consistently satisfies the sufficient descent property, regardless of the line search employed.
-To demonstrate the effectiveness of the proposed approach, we present numerical results that validate its performance and show its superior convergence compared to existing methods.
+This new method consistently satisfies the sufficient descent property, regardless of the line search used.
+To demonstrate the effectiveness of the proposed approach, we present numerical results that validate its performance and show its superior convergence compared to the modified spectral conjugate gradient method of Faramarzi and Amini [26] and the scaled conjugate gradient method of Mard and Fakhari [28].
 Furthermore, we explore its application in compressed sensing, emphasizing its practical relevance and significant advantages.
@@
 Let $\mathbb{R}^{m}$ be the Euclidean space and let $\{u_{n}\}$ and $\{v_{n}\}$ be sequences in $\mathbb{R}^{m}$.
-Let $\{a_{n}\}$ and $\{b_{n}\}$ be sequences in $(0,\infty)$ and $u_{0}\in\mathbb{R}^{m}$ and $b_{0},\sigma\in(0,\infty)$ satisfying
+Let $b_{0},\sigma\in(0,\infty)$ and $u_{0}\in\mathbb{R}^{m}$ and let $\{a_{n}\}$ and $\{b_{n}\}$ be sequences in $(0,\infty)$ satisfying the following conditions:
   1. (C1)
@@
 $-\langle v_{n+1},u_{n+1}\rangle+\sigma b_{n}|\langle u_{n},v_{n}\rangle|\geq a_{n+1}\|v_{n+1}\|^{2}.$
-Squaring both side and applying Lemma 1 with $c=-\langle v_{n+1},u_{n+1}\rangle>0$ (from (C2)) and $d=b_{n}|\langle v_{n},u_{n}\rangle|>0$ gives
+Squaring both sides and applying Lemma 1 with $c=-\langle v_{n+1},u_{n+1}\rangle>0$ (from (C2)) and $d=b_{n}|\langle v_{n},u_{n}\rangle|>0$ gives
 $\langle v_{n+1},u_{n+1}\rangle^{2}+b_{n}^{2}\langle v_{n},u_{n}\rangle^{2}\geq\frac{a_{n+1}^{2}}{1+\sigma^{2}}\|v_{n+1}\|^{4}.$
@@
 which implies that
-$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}$ $=\frac{1}{\|u_{n+1}\|^{2}}\left[\langle v_{n+1},u_{n+1}\rangle^{2}+\frac{\|u_{n+1}\|^{2}}{\|u_{n}\|^{2}}\langle v_{n},u_{n}\rangle^{2}\right]$
+$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}$
+$=\frac{1}{\|u_{n+1}\|^{2}}\left[\langle v_{n+1},u_{n+1}\rangle^{2}+\frac{\|u_{n+1}\|^{2}}{\|u_{n}\|^{2}}\langle v_{n},u_{n}\rangle^{2}\right]$
 $\geq\frac{1}{\|u_{n+1}\|^{2}}\left[\frac{a}{1+\sigma^{2}}\|v_{n+1}\|^{4}+\left(\frac{\|u_{n+1}\|^{2}}{\|u_{n}\|^{2}}-b_{n}^{2}\right)\langle v_{n},u_{n}\rangle^{2}\right].$ (15)
@@
 $\sum_{n=0}^{\infty}\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}<\infty.$
-Remark 5.
 From Lemma 4, we have
@@
 3 A modified descent Dai-Liao spectral conjugate gradient method and its convergence analysis
-In this section, inspired by the concepts from quasi-Newton methods, we introduce an effective modified spectral conjugate gradient (CG) method.
-Motivated by the theoretical properties provided by (7) and the Dai-Liao conjugate gradient parameter (5), we propose the following modified Dai-Liao conjugate parameter:
+In this section, we introduce a modified spectral conjugate gradient (CG) method, inspired by ideas from quasi-Newton methods.
+Using the theoretical properties provided by (7) and the Dai-Liao conjugate gradient parameter (5), we propose the following modified Dai-Liao conjugate parameter:
 $\beta_{n}=\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-t\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0},$ (17)
@@
 Following the methodology in [11], we propose the following formula for the parameter $t$ in the modified Dai-Liao conjugate parameter (17):
-$t_{n}^{p,q}=p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\text{ for all }n\in\mathbb{N}_{0},$ (18)
+$t_{n}^{p,q}=p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\text{ for all }n\in\mathbb{N}_{0}$ (18)
 with $p\in\left(\frac{1}{4},\infty\right)$ and $q\in\left(-\infty,\frac{1}{4}\right)$. Using this parameter $t_{n}^{p,q}$, we define the following modified descent Dai-Liao (MDDL) conjugate parameter:
@@
 where $\{\theta_{n}\}$ is a sequence in $(0,\infty)$.
-Pproposition 6.
-Let $f:\mathbb{R}^{m}\to\mathbb{R}$ be a differentiable function and let $\{d_{n}\}$ be a sequence of spectral CG search directions generated by (20). Then we have the following:
+Proposition 5.
+Let $f:\mathbb{R}^{m}\to\mathbb{R}$ be a differentiable function and let $\{d_{n}\}$ be a sequence of spectral CG search directions generated by (20). Then, we have the following:
   • (a)
     For all $n\in\mathbb{N}_{0},$ we have
-    $\langle\nabla f(x_{n}),d_{n}\rangle\leq-\kappa_{n}\|\nabla f(x_{n})\|^{2}$ (21)
+    $\langle\nabla f(x_{n}),d_{n}\rangle\leq-\kappa_{n}\|\nabla f(x_{n})\|^{2},$ (21)
     where
-    $\kappa_{n}=\left\{\begin{array}[]{ll}1&\text{if }n=0,\\ \theta_{n}-\left(\frac{1}{4p}+|q|\right)&\text{if }n\geq 1.\end{array}\right.$
+    $\kappa_{n}=\left\{\begin{array}[]{ll}1,&\text{if }n=0,\\ \theta_{n}-\left(\frac{1}{4p}+|q|\right),&\text{if }n\geq 1.\end{array}\right.$
   • (b)
@@
 From (18), (19) and (20), we get
-$\langle\nabla f(x_{n+1}),d_{n+1}\rangle$
-$=\langle\nabla f(x_{n+1}),-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}^{MDDL}d_{n}\rangle$
+$\langle\nabla f(x_{n+1}),d_{n+1}\rangle$ $=\langle\nabla f(x_{n+1}),-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}^{MDDL}d_{n}\rangle$
 $=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
@@
 $=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
-$\quad-p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}\frac{\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}+q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\frac{\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}.$
+$\quad-p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}\frac{\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}$
+$\quad+q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\frac{\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}.$
 Since $s_{n}=\alpha_{n}d_{n}$, we have
-$\langle\nabla f(x_{n+1}),d_{n+1}$
-$=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}-p\frac{\|z_{n}\|^{2}\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle}$
-$\quad+q\frac{\langle d_{n},z_{n}\rangle}{\|d_{n}\|^{2}}\frac{\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle}$
-$=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}-p\frac{\|z_{n}\|^{2}\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle^{2}}$
-$\quad+q\frac{\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\|d_{n}\|^{2}}.$ (22)
+$\langle\nabla f(x_{n+1}),d_{n+1}\rangle$ $=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
+$\quad-p\frac{\|z_{n}\|^{2}\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle}+q\frac{\langle d_{n},z_{n}\rangle}{\|d_{n}\|^{2}}\frac{\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle}$
+$=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
+$\quad-p\frac{\|z_{n}\|^{2}\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle^{2}}+q\frac{\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\|d_{n}\|^{2}}.$ (22)
 From Lemma 2 with $u_{n}=\frac{1}{\sqrt{2p}}\nabla f(x_{n+1})$ and $v_{n}=\frac{\sqrt{2p}\langle\nabla f(x_{n+1}),d_{n}\rangle z_{n}}{\langle d_{n},z_{n}\rangle},$ we obtain
@@
 Since $B_{n+1}$ must satisfies the modified secant condition $B_{n+1}s_{n}=z_{n}$, we have
-${\theta_{n+1}}$
-$=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}(\langle s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle z_{n},d_{n}\rangle)$
-$=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}\left(\langle s_{n},\nabla f(x_{n+1})\rangle+\left(\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right)\langle z_{n},d_{n}\rangle\right)$
-$=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}\left(\langle s_{n},\nabla f(x_{n+1})\rangle+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle z_{n},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle\langle z_{n},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right)$
-$=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}\left(\langle s_{n},\nabla f(x_{n+1})\rangle+\langle\nabla f(x_{n+1}),z_{n}\rangle-t_{n}^{p,q}\langle\nabla f(x_{n+1}),s_{n}\rangle\right)$
-$=1-\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}\left({t_{n}^{p,q}\langle s_{n},\nabla f(x_{n+1})\rangle}-\langle s_{n},\nabla f(x_{n+1})\rangle\right)$ (25)
+${\theta_{n+1}}$ $=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}(\langle s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle z_{n},d_{n}\rangle)$
+$=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}\left(\langle s_{n},\nabla f(x_{n+1})\rangle+\left(\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right.\right.$
+$\quad\left.\left.-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right)\langle z_{n},d_{n}\rangle\right)$
+$=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}\left(\langle s_{n},\nabla f(x_{n+1})\rangle+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle z_{n},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right.$
+$\quad\left.-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle\langle z_{n},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right)$
+$=\frac{\langle s_{n},\nabla f(x_{n+1})\rangle+\langle\nabla f(x_{n+1}),z_{n}\rangle-t_{n}^{p,q}\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle z_{n},\nabla f(x_{n+1})\rangle}$
+$=1-\frac{{t_{n}^{p,q}\langle s_{n},\nabla f(x_{n+1})\rangle}-\langle s_{n},\nabla f(x_{n+1})\rangle}{\langle z_{n},\nabla f(x_{n+1})\rangle}$ (25)
 $=1-(t_{n}^{p,q}-1)\frac{\langle s_{n},\nabla f(x_{n+1})\rangle}{\langle z_{n},\nabla f(x_{n+1})\rangle}.$ (26)
@@
 ${\theta_{n+1}}=1-\frac{t_{n}^{p,q}\langle s_{n},\nabla f(x_{n+1})\rangle}{\langle z_{n},\nabla f(x_{n+1})\rangle}.$ (27)
-We denote the $\theta_{n+1}$ obtained in equation (26) by $\theta_{n+1}^{N+}$ and the $\theta_{n+1}$ obtained in equation (27) by $\theta_{n+1}^{N-}$.
+We denote the $\theta_{n+1}$ obtained in equation (26) by $\theta_{n+1}^{R}$ and the $\theta_{n+1}$ obtained in equation (27) by $\theta_{n+1}^{N}$.
 To obtain the sufficient descent condition and the bounded property of the spectral parameters, we propose the following spectral parameters:
@@
 Now, we propose the following modified descent Dai-Liao spectral conjugate gradient method (MDDLSCG):
-Input: Let $x_{0}\in\mathbb{R}^{n}$ and let $\epsilon,\delta,\sigma,\eta,\tau,r,\nu\in(0,\infty)$ such that $10<\delta<\sigma<$ 1.
+Input: Let $x_{0}\in\mathbb{R}^{n}$ and let $\epsilon,\delta,\sigma,\eta,\tau,r,\nu\in(0,\infty)$ such that
+     $10<\delta<\sigma<$ 1.
 Step 0: Set $d_{0}=-\nabla f(x_{0})$.
 Step 1: If $\|\nabla f(x_{n})\|_{\infty}<\epsilon$, stop.
-Step 2: Determine the step length $\alpha_{n}$ satisfying the strong Wolfe conditions (8) and (9).
+Step 2: Determine the step length $\alpha_{n}$ satisfying the strong Wolfe
+    conditions (8) and (9).
 Step 3: Set $x_{n+1}=x_{n}+\alpha_{n}d_{n}$.
@@
 $d_{n+1}=-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}d_{n}\text{ for all }n\in\mathbb{N}_{0},$
-where CG parameter $\beta_{n}^{MDDL}$ is given by (19) and spectral parameter is given by (28) or (29).
+      where CG parameter $\beta_{n}^{MDDL}$ is given by (19) and spectral
+    parameter is given by (28) or (29).
 Step 5: Set $n=n+1$ and return to Step 1.
@@
 $\langle d_{n},y_{n}\rangle\geq-(1-\sigma)\langle\nabla f(x_{n}),d_{n}\rangle\geq(1-\sigma)\eta\|\nabla f(x_{n})\|^{2}>0,$ (32)
-which implies, from (7) that $h_{n}=\nu$ and hence
-$z_{n}=y_{n}+\nu\|\nabla f(x_{n})\|^{r}s_{n},$ (33)
-where $r,\nu\in(0,\infty)$.
+which, from (7), implies that $h_{n}=\nu$ and hence
+$z_{n}=y_{n}+\nu\|\nabla f(x_{n})\|^{r}s_{n}.$ (33)
 Now, we analyze the global convergence of the MDDLSCG Algorithm.
@@
 Lemma 8.
-Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a function and let $\{x_{n}\}$ be sequence generated by Algorithm 1 such that Assumptions (A1) and (A2) hold.
+Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a function and let $\{x_{n}\}$ be the sequence generated by Algorithm 1 such that Assumptions (A1) and (A2) hold.
 Suppose that $\theta_{n}>\frac{1}{4p}+|q|$ for all $n\in\mathbb{N}$.
@@
 Also from strong Wolfe condition (9), we have $|\langle\nabla f(x_{n+1}),d_{n}\rangle|\leq-\sigma\langle\nabla f(x_{n}),d_{n}\rangle$.
-Then, clearly all the conditions (C1), (C2) and (C3) of Lemma 3 is satisfied.
+Then, clearly all the conditions (C1), (C2) and (C3) of Lemma 3 are satisfied.
 Thus, we have
@@
 $\frac{\langle\nabla f(x_{n+1}),d_{n+1}\rangle^{2}}{\|d_{n+1}\|^{2}}+\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}\geq b_{1}\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}.$ (36)
-From (36) and Lemma 4, we get $\sum_{n=1}^{\infty}\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}<\infty.$
+From (36) and Lemma 4, we get that $\sum_{n=1}^{\infty}\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}<\infty.$
 Theorem 9.
-Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a function and let $\{x_{n}\}$ be sequence generated by Algorithm 1 such that Assumptions (A1) and (A2) hold. Then $\liminf_{n\to\infty}\|\nabla f(x_{n})\|=0.$
+Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a function and let $\{x_{n}\}$ be the sequence generated by Algorithm 1 such that Assumptions (A1) and (A2) hold.
+Then $\liminf_{n\to\infty}\|\nabla f(x_{n})\|=0.$
 Proof.
@@
 From (32), (33) and (37), we obtain
-$\langle d_{n},z_{n}\rangle=\langle d_{n},y_{n}\rangle+\nu\|\nabla f(x_{n})\|^{r}\alpha_{n}\|d_{n}\|^{2}\geq\nu\|\nabla f(x_{n})\|^{r}\alpha_{n}\|d_{n}\|^{2}\geq\nu\xi^{r}\alpha_{n}\|d_{n}\|^{2}.$ (38)
+$\langle d_{n},z_{n}\rangle$ $=\langle d_{n},y_{n}\rangle+\nu\|\nabla f(x_{n})\|^{r}\alpha_{n}\|d_{n}\|^{2}$
+$\geq\nu\|\nabla f(x_{n})\|^{r}\alpha_{n}\|d_{n}\|^{2}\geq\nu\xi^{r}\alpha_{n}\|d_{n}\|^{2}.$ (38)
 From (33), (10) and ((ii)), we have
@@
 From the Lipschitz continuity of the $\nabla f$ and (39), we obtain
-$t_{n}^{p,q}\leq\left|p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\right|\leq p\frac{(L+\nu\gamma^{r})^{2}\|s_{n}\|^{2}}{\mu\|s_{n}\|^{2}}+|q|\frac{(L+\nu\gamma^{r})\|s_{n}\|^{2}}{\|s_{n}\|^{2}}=p\frac{L^{2}}{\mu}+L|q|.$
+$t_{n}^{p,q}$ $\leq\left|p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\right|$
+$\leq p\frac{(L+\nu\gamma^{r})^{2}\|s_{n}\|^{2}}{\mu\|s_{n}\|^{2}}+|q|\frac{(L+\nu\gamma^{r})\|s_{n}\|^{2}}{\|s_{n}\|^{2}}=p\frac{L^{2}}{\mu}+L|q|.$
 Thus, the relations (19), (10), ((ii)), (38), (39), along with Cauchy-Schwarz inequality, conclude
-$\|\beta_{n}^{MDDL}\|$ $=\left\|\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
+$\|\beta_{n}^{MDDL}\|$
+$=\left\|\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
 $=\left\|\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-\left(p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\right)\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
 $=\left\|\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}+q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
-$=\left\|\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|+p\left\|\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|+|q|\left\|\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
+$=\left\|\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|+p\left\|\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
+$\quad+|q|\left\|\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right\|$
 $=\frac{\|\nabla f(x_{n+1})\|\|z_{n}\|}{\langle d_{n},z_{n}\rangle}+p\frac{\|z_{n}\|^{2}\|\nabla f(x_{n+1})\|\|d_{n}\|}{\langle d_{n},z_{n}\rangle^{2}}+|q|\frac{\|\nabla f(x_{n+1})\|}{\|d_{n}\|}$ (40)
@@
 $\frac{1}{\|d_{n+1}\|^{2}}\geq\frac{1}{E^{2}}\textrm{ for all }n\in\mathbb{N}_{0},$
-where $E=\tau\gamma+\frac{\gamma(L+\nu\gamma^{r})}{\nu\xi^{r}}+\frac{p\gamma(L+\nu\gamma^{r})^{2}}{\nu^{2}\xi^{2r}}+|q|\gamma$. This implies that
+where $E=\tau\gamma+\frac{\gamma(L+\nu\gamma^{r})}{\nu\xi^{r}}+\frac{p\gamma(L+\nu\gamma^{r})^{2}}{\nu^{2}\xi^{2r}}+|q|\gamma$. Thus,
 $\sum_{n=0}^{\infty}\frac{1}{\|d_{n}\|^{2}}=\infty.$ (42)
@@
 4 Numerical Experiments
-In this section, we discuss the numerical performance of our proposed method (MDDLSCG), comparing it with Algorithm 3.1 in [26] (denoted by MSCG) and Scaled conjugate gradient method (denoted by ScCGM) in [29].
-All numerical experiments were conducted in MATLAB R2024b on an HP workstation with Intel Core i7-13700 processor, 32 GB RAM and Window 11.
-During the numerical experiments, we set the parameters $\delta=0.01$, $\sigma=0.1$ for all the algorithms, $\eta=0.001$, $\tau=10$, $r=1$ and $\nu=0.001$ for algorithms MSCG and MDDLSCG and $p=0.4$ and $q=0.2$ for the algorithm MDDLSCG.
-During the implementation, the iterations are terminated, when it meets the stopping criterion $E_{n}=\|\nabla f(x_{n})\|_{\infty}<10^{-15}$.
-For the numerical experiment, we consider the test problem of finding the minima of Beale function.
-The Beale function is multimodal with sharp peaks at the corners of the input domain, which is defined from the general Beale function as follows:
+In this section, we discuss the numerical performance of our proposed method (MDDLSCG), comparing it with Algorithm 3.1 in [26] (denoted by MSCG) and Scaled conjugate gradient method (denoted by ScCG) in [28].
+All numerical experiments were conducted in MATLAB R2024b on an HP workstation with Intel Core i7-13700 processor, 32 GB RAM and Windows 11.
+For the numerical experiment, we consider two test problems: (i) finding the minima of the Beale function, and (ii) finding the minima of the Almost Perturbed Quadratic (APQ) function:
+Beale Test Function: The Beale function is multimodal with sharp peaks at the corners of the input domain, which is defined from the general Beale function as follows:
 $f(x,y)=\left(1.5-x+xy\right)^{2}+\left(2.25-x+xy^{2}\right)^{2}+\left(2.625-x+xy^{3}\right)^{2}\text{ for all }(x,y)\in\mathbb{R}^{2}.$
 The global minimum is located at $(x,y)=(3,0.5)$, where $f(x,y)=0$.
-The numerical results of all the algorithms (MDDLSCG, MSCG and ScCG) are summarized in Table 1, while their performance is illustrated in Figure LABEL:bealee.
-In Figure LABEL:bealee (a), the iteration paths are visualized on the contour plot of the Beale function, showing how each method converges to the minimizer.
-Figure LABEL:bealee (b) depicts the convergence rate by plotting the norm of the gradient, $\|\nabla f(x_{n})\|_{\infty}$ versus the number of iterations on a logarithmic scale.
-From the data presented, we can clearly observe that our method, MDDLSCG, outperforms the MSCG algorithm and ScCG in all tested scenarios.
-This superiority is evidenced by the metrics shown in the table and figures, which highlight the efficiency and effectiveness of MDDLSCG compared to MSCG and ScCG.
-Method itr Tcpu $E_{n}$
-MDDLSCG 21 5.957900e-03 3.580469e-15
-MSCG 30 1.783450e-02 4.662937e-15
-ScCG 139 9.749100e-03 8.840151e-15
-Table 1: Comparison table for MDDLSCG, MSCG and ScCG for Beale problems
+During the numerical experiments, we set the parameters $\delta=0.01$, $\sigma\in\{0.1,0.2,0.4,0.6,0.8,0.9\}$ for all the algorithms, $\eta=0.001$, $\tau=10$, $r=1$ and $\nu=0.001$ for algorithms MSCG and MDDLSCG and $p=0.4$ and $q=0.2$ for the algorithm MDDLSCG.
+We take the initial point $x_{0}=(1,0.8)$.
+During the implementation, the iterations are terminated, when it meets the stopping criterion $E_{n}=\|\nabla f(x_{n})\|_{\infty}<10^{-15}$.
+The numerical results of all the algorithms MDDLSCG, MSCG and ScCG for the Beale function are summarized in Table 1, while their performance is illustrated in Figure LABEL:bealefig1 and LABEL:bealefig2.
+In Figures LABEL:bealefig1 (a), (c), (e), the iteration paths are visualized on the contour plot of the Beale function, showing how each method converges to the minimizer for $\sigma=0.1,0.2,0.6$, respectively.
+Figures LABEL:bealefig1 (b), (d), (f) depict the convergence rate by plotting the norm of the gradient, $\|\nabla f(x_{n})\|_{\infty}$ versus the number of iterations on a logarithmic scale for $\sigma=0.1,0.2,0.6$, respectively.
+The graph between different values of sigma and number of iterations is given in Figure LABEL:bealefig2 (a) and the graph between different values of sigma and execution time is in Figure LABEL:bealefig2 (b).
+From the data presented, it is evident that our proposed method, MDDLSCG, outperforms both MSCG and ScCG in all tested scenarios.
+As shown in Table 1, Figure LABEL:bealefig1 and LABEL:bealefig2 (b), MDDLSCG reaches the stopping criteria in fewer iterations and requires less execution time.
+Additionally, From Figure LABEL:bealefig2 (a), we observe that these methods are significantly sensitive to variations in the parameter $\sigma$, which affects their performance.
+$\sigma$ MDDLSCG MSCG ScCG itr Tcpu $E_{n}$ itr Tcpu $E_{n}$ itr Tcpu $E_{n}$ 0.1 37 1.86e-02 9.49e-15 63 3.14e-02 2.63e-16 200 5.40e-02 5.27e-16 0.2 36 1.75e-02 3.16e-15 63 2.76e-02 9.07e-15 136 3.99e-02 4.51e-15 0.4 53 2.44e-02 9.40e-15 73 2.65e-02 8.75e-15 126 4.73e-02 2.87e-15 0.6 40 2.36e-02 7.91e-16 72 2.78e-02 5.74e-15 95 4.94e-02 9.32e-15 0.8 67 3.24e-02 5.32e-15 111 2.86e-02 4.48e-15 201 4.92e-02 6.91e-15 0.9 64 3.28e-02 8.16e-15 101 3.48e-02 8.74e-15 184 4.37e-02 1.08e-15
+Table 1: Comparison table for MDDLSCG, MSCG and ScCG for Beale function
+Almost Perturbed Quadratic function: Consider the Almost Perturbed Quadratic (APQ) function $f:\mathbb{R}^{m}\to\mathbb{R}$ defined by
+$f(\mathbf{x})=\frac{1}{2}\mathbf{x}^{T}A\mathbf{x}+\mathbf{b}^{T}\mathbf{x}$
+where $A\in\mathbb{R}^{m\times m}$ is a symmetric positive definite matrix ans $b\in\mathbb{R}^{m}$ is a vector of perturbation.
+This function has a global minimum at $x=-A^{-1}b$.
+Note that the gradient of $f$ is given by
+$\nabla f(x)=Ax+b\text{ for all }x\in\mathbb{R}^{m}.$
+For the numerical experiments, we have considered three cases, each characterized by dimensions of1000100010001000, $100$ and $25$.
+To the cases with dimensions $1000$ and $100$, the matrix $A$ and vector $b$ are randomly generated using MATLAB’s built-in function rand and the initial point is randomly generated using MATLAB’s rand function.
+For the case with dimension $25$, we take the initial point
+$x_{0}$ $=(x_{0,1},\ldots,x_{0,i},\ldots,x_{0,25}),\text{ where }x_{0,i}=1\text{ for all }i=1,2,\ldots,25\text{ and}$
+$a$ $=(2,4,3,6,4,12,8,10,5,8,6,11,4,8,3,6,2,12,6,8,4,10,5,2,6),$
+$b$ =(4,52,18,60,24,72,72,50,55,80,78,22,12=(-4,-52,-18,-60,-24,-72,-72,-50,-55,-80,-78,-22,-12,= ( - 4 , - 52 , - 18 , - 60 , - 24 , - 72 , - 72 , - 50 , - 55 , - 80 , - 78 , - 22 , - 12 ,
+80,27,42,24,120,30,120,24,100,60,20,66)$\quad$-80,-27,-42,-24,-120,-30,-120,-24,-100,-60,-20,-66),- 80 , - 27 , - 42 , - 24 , - 120 , - 30 , - 120 , - 24 , - 100 , - 60 , - 20 , - 66 ) ,
+$A$ $=\text{diag}(a).$
+For all the cases, the parameters are set as follows: $\delta=0.01$ and $\sigma=0.1$ for all algorithms, while for MSCG and MDDLSCG, we use $\eta=0.001$, $\tau=10$, $r=1$, and $\nu=0.001$.
+Additionally, for the MDDLSCG algorithm, we set $p=0.4$ and $q=0.2$.
+During the implementation, the iterations are terminated when the stopping criterion $E_{n}=\|\nabla f(x_{n})\|_{\infty}<10^{-6}$ is satisfied, where $x_{n}=(x_{n,1},\ldots,x_{n,25})$.
+The performance of the algorithms MDDLSCG, MSCG, and ScCG for the APQ function is summarized in Table 2 for all three cases.
+To provide a clearer visualization of the experiment, we have plotted the graph of the number of iterations versus $E_{n}$ in Figure LABEL:APQgraph.
+From Table 2 and Figure LABEL:APQgraph, it is evident that our proposed method, MDDLSCG, requires fewer iterations and less computational time to satisfy the desired stopping criterion across all cases.
+Figure LABEL:APQtrack illustrates how the algorithms MDDLSCG, MSCG, and ScCG approximate the desired solution $x^{*}=-A^{-1}b$ at the 5th and 10th iterations for the case where the problem dimension is $25$.
+Furthermore, Figure LABEL:APQprogress depicts the convergence behavior of the algorithms as the number of iterations increases up to the 50th iteration for the same dimension.
+It is observed that MDDLSCG demonstrates superior convergence characteristics compared to MSCG and ScCG.
+Dimension Method itr Tcpu $Er_{n}$
+25 MDDLSCG 52 3.088055e-02 5.691882e-07
+MSCG 59 3.947327e-02 6.931169e-07
+ScCG 65 4.730704e-02 4.514777e-07
+100 MDDLSCG 102 5.552326e-02 8.162825e-07
+MSCG 121 6.512598e-02 9.411660e-07
+ScCG 134 6.823947e-02 8.609522e-07
+1000 MDDLSCG 132 3.481332e-02 9.291368e-07
+MSCG 146 3.522395e-02 9.484382e-07
+ScCG 171 4.053286e-02 8.747078e-07
+Table 2: Comparison table of MDDLSCG, MSCG and ScCG for APQ function
 5 Application in compressed sensing
@@
 However, solving BPD is challenging due to the non-smoothness of the $\ell_{1}$-penalty term.
-To address this, Zhu et al.
+To address this issue, Zhu et al.
 [33] proposed a relaxation of BPD using Nesterov’s smoothing technique [34], resulting in the following formulation:
@@
 where $x_{i}$ represents the $i$-th component of the vector $x$, and $\psi_{\lambda}$ is defined as:
-$\psi_{\lambda}(|x_{i}|)=\max_{\gamma\in[0,1]}\left\{\gamma|x_{i}|-\frac{1}{2}\lambda\gamma^{2}\right\}$
+$\psi_{\lambda}(|x_{i}|)=\max_{\gamma\in[0,1]}\left\{\gamma|x_{i}|-\frac{1}{2}\lambda\gamma^{2}\right\},$
 which simplifies to:
@@
 $\nabla f(x)=\lambda\nabla f_{\lambda}(x)+A^{T}(Ax-b).$
-For the numerical experiments, we aim to reconstruct a sparse signal of length $n$ from $m$-length observations ($m\ll n$).
-The parameters for the MDDLSCG, MSCG and ScCG algorithms are set as follows: $\mu=\max\{2^{-7},0.001\|A^{T}b\|_{\infty}\}$, $\lambda=\min\{0.001,0.048\|A^{T}b\|_{\infty}\}$, $\sigma=0.1$, $\delta=0.01$, $\eta=0.001$, $\tau=10$, $r=1$ and $\nu=0.001$.
+For the numerical experiments, we aim to reconstruct a sparse signal of length $n$ from $m$-length observations ($m\ll n$). The parameters for the MDDLSCG, MSCG and ScCG algorithms are set as follows:
+$\mu=\max\{2^{-7},0.001\|A^{T}b\|_{\infty}\}$, $\lambda=\min\{0.001,0.048\|A^{T}b\|_{\infty}\}$, $\sigma=0.1$, $\delta=0.01$, $\eta=0.001$, $\tau=10$, $r=1$ and $\nu=0.001$.
 […]
 $\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(x_{i}-x_{i}^{*})^{2}\leq 10^{-5}.$
-To evaluate the restoration performance qualitatively, we report the relative error (RelErr) of the recovered signal, defined as:
+To evaluate the restoration performance qualitatively, we report the relative error (RelErr) of the recovered signal, which is defined as follows:
 $\text{RelErr}=\frac{\|x^{*}-x\|}{\|x\|}.$
@@
 The graph depicting the relationship between relative error and the number of iterations is shown in Figures LABEL:stpp512(a) and LABEL:stopp(a), while the graph illustrating MSE versus the number of iterations is presented in Figures LABEL:stpp512(b) and LABEL:stopp(b).
-Additionally, the recovery of the signal is visualized in Figures 4 and 6.
+Additionally, the recovery of the signal is visualized in Figures 11 and 13.
 From Table 3, Figures LABEL:stpp512, and LABEL:stopp, it is clear that the MDDLSCG algorithm surpasses the MSCG and ScCG algorithms in performance across all cases.
@@
 Figure 13: Signal recovery for data $(m,n,k)=(256,1024,32)$
-Dimensions $(m,n,k)$ Method Itr Tcpu MSE RelErr
-(128,512,16) MDDLSCG 272 1.186382e-01 2.033486e-06 7.487872e-03
-MSCG 493 1.606987e-01 2.033495e-06 7.488087e-03
-ScCG 3653 1.135658e+00 2.033588e-06 7.712389e-03
-(256,1024,32) MDDLSCG 291 2.163503e-01 2.128598e-06 9.721181e-03
-MSCG 1405 8.339790e-01 2.128516e-06 9.720993e-03
-ScCG 3002 1.746555e+00 2.229608e-04 9.949161e-02
+Dim$(m,n,k)$ Method Itr Tcpu MSE RelErr (128,512,16) MDDLSCG 272 1.186382e-01 2.033486e-06 7.487872e-03 MSCG 493 1.606987e-01 2.033495e-06 7.488087e-03 ScCG 3653 1.135658e+00 2.033588e-06 7.712389e-03 (256,1024,32) MDDLSCG 291 2.163503e-01 2.128598e-06 9.721181e-03 MSCG 1405 8.339790e-01 2.128516e-06 9.720993e-03 ScCG 3002 1.746555e+00 2.229608e-04 9.949161e-02
 Table 3: Comparison table for methods MDDLSCG, MSCG and ScCG
 6 Conclusions
-We presented a modified descent Dai-Liao spectral conjugate gradient (MDDLSCG) method that incorporates a new secant condition and quasi-Newton direction.
+We investigated a modified descent Dai-Liao spectral conjugate gradient (MDDLSCG) method that incorporates a new secant condition and quasi-Newton direction.
 The updated method introduces spectral parameters ensuring a sufficient descent property independent of line search techniques.
 Theoretical analysis establishes the global convergence of the method for general nonlinear functions under standard assumptions.
-Numerical experiments and an application in signal processing demonstrate the improved performance and practical effectiveness of the proposed algorithm, highlighting its potential in optimization and related applications.
+Numerical experiments and an application in signal processing demonstrate the improved performance and practical effectiveness of the proposed algorithm.
+Additionally, we present numerical results that validate its performance and show its superior convergence compared to the MSCG method [26] and the ScCG method [28], highlighting its potential in optimization and related applications.
+Declaration of competing interest
+The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+Acknowledgments
+The authors are grateful to the editor and to the reviewers for their remarks and advice which improved our manuscript.
+Funding
+The authors declare that no funds, grants, or other support were received during the preparation of this manuscript.

### v3

@@
-Modified Dai-Liao Spectral Conjugate Gradient Method with Application to Signal Processing
+A Modified Dai-Liao Spectral Conjugate Gradient Method with an Application to Signal Processing
 Abstract
-In this article, we present a modified variant of the Dai-Liao spectral conjugate gradient method, developed through an analysis of eigenvalues and inspired by a modified secant condition.
-We show that the proposed method is globally convergent for general nonlinear functions under standard assumptions.
+We propose and study a variant of the Dai-Liao spectral conjugate gradient method, developed through an analysis of eigenvalues and inspired by a modified secant condition.
+We show that our proposed method is globally convergent for general nonlinear functions under standard assumptions.
 By incorporating the new secant condition and a quasi-Newton direction, we introduce updated spectral parameters.
@@
 Numerical experiments show that the proposed algorithm performs better than several existing methods in terms of convergence speed and computational efficiency.
-Its effectiveness is further demonstrated through an application in signal processing.
+Its effectiveness is further demonstrated through an application to signal processing.
 keywords:
-Spectral Conjugate Gradient Method , Dai-Liao Method , Optimization , Signal Processing
-MSC:
-[2020] 65K05 , 90C30 , 49M37
+Dai-Liao Method , Optimization , Signal Processing , Spectral Conjugate Gradient Method
+2020 MSC:
+65K05 , 90C30 , 49M37
 […]
@@
 […]
-organization=Department of Mathematics, The Technion- Israel Institute of Technology,city=Haifa, country=Israel
+organization=Department of Mathematics, The Technion – Israel Institute of Technology,city=Haifa, country=Israel
 […]
 organization=Department of Applied Mathematics and Scientific Computing, Indian Institute of Technology Roorkee,country=India
+[…]
+organization=Department of Mathematics, The Technion- Israel Institute of Technology,city=Haifa, country=Israel
 1 Introduction
-[…]
-find $\text{find }x^{*}\in\mathbb{R}^{m}\text{ such that }f(x^{*})=\min_{x\in\mathbb{R}^{m}}f(x).$ (1)
-We denote the set of minimizers of $f$ as $\Omega_{\text{min}}(f,\mathbb{R}^{m})$.
-The general iterative rule for solving (1) starts from an initial point $x_{0}\in\mathbb{R}^{m}$ and generates a sequence $\{x_{n}\}$ in $\mathbb{R}^{m}$ as follows:
+Let $\mathbb{R}^{m}$ denote the Euclidean space equipped with the usual inner product $\langle\cdot,\cdot\rangle$ and its associated norm $\|\cdot\|$.
+Let $f:\mathbb{R}^{m}\to\mathbb{R}$.
+Our goal is to solve the following minimization problem:
+$\text{find }x^{*}\in\mathbb{R}^{m}\text{ such that }f(x^{*})=\min_{x\in\mathbb{R}^{m}}f(x).$ (1)
+We denote the set of minimizers of $f$ by $\Omega_{\text{min}}(f,\mathbb{R}^{m})$.
+A common approach to find a solution to (1) is the following iterative method: Let $x_{0}\in\mathbb{R}^{m}$ be an initial point.
+Given the current iterate $x_{n}$, the next iterate $x_{n+1}$ is computed by
 $x_{n+1}=x_{n}+\alpha_{n}d_{n}\text{ for all }n\in\mathbb{N}_{0}=\mathbb{N}\cup\{0\},$ (2)
-where the stepsize $\alpha_{n}$ is a positive real parameter determined via exact or inexact line search, $x_{n}$ is the previous iterative point, $x_{n+1}$ is the current point and $d_{n}$ is a suitable search direction.
-The class of algorithms that follow the form (2) is known as line search algorithms.
-For such algorithms, only the search direction $d_{n}\in\mathbb{R}^{m}$ and the stepsize $\alpha_{n}\in(0,\infty)$ are required.
-An appropriate descent direction $d_{n}$ must be chosen such that it satisfies the descent condition:
+where $\alpha_{n}>0$ is the stepsize determined by exact or inexact line search and $d_{n}$ is a suitable search direction.
+Methods that follow the update rule (2) are referred to as line search methods.
+These methods require only the search direction $d_{n}\in\mathbb{R}^{m}$ and the stepsize $\alpha_{n}\in(0,\infty)$.
+An appropriate descent direction $d_{n}$ must be chosen so that it satisfies the descent condition:
 $\langle\nabla f(x_{n}),d_{n}\rangle<0\text{ for all }n\in\mathbb{N}_{0}.$ (3)
-A common choice for a descent direction is $d_{n}=-\nabla f(x_{n})$, which reduces the general line search iteration (2) to the gradient descent (GD) method:
+A common choice for a descent direction is $d_{n}=-\nabla f(x_{n})$, which simplifies the general line search method (2) to the familiar gradient descent (GD) method:
 $x_{n+1}=x_{n}-\alpha_{n}\nabla f(x_{n})\text{ for all }n\in\mathbb{N}_{0}.$
@@
 Conjugate gradient (CG) methods are among the most effective algorithms for solving large-scale unconstrained optimization problems due to their low storage requirements.
-These methods require only first-order derivatives, making them highly suitable for large-scale applications.
-The first form of the CG method was introduced by Hestenes and Stiefel [1] in 1952 for solving linear systems of equations.
-Later, in 1964, Fletcher and Reeves [2] modified the original idea and developed the first nonlinear CG method for unconstrained optimization problems.
-Due to the strong theoretical foundation and excellent numerical performance of these algorithms, many variants of the CG method have been proposed (see, e.g., [3, 4, 5, 6]).
-The successive approximations generated by the CG algorithms have the form (2), where $\alpha_{n}\in(0,\infty)$ is the step length determined through a line search along the CG direction $d_{n}$.
-The CG directions $d_{n}$ is defined as:
+These methods require only first-order derivatives, makes them highly suitable for large-scale applications.
+The conjugate gradient (CG) method was originally proposed by Hestenes and Stiefel [Hestenes1952] for the solution of linear systems.
+Later, Fletcher and Reeves [Fletcher] extended this concept and formulated the first nonlinear CG method for unconstrained optimization.
+Due to their strong theoretical foundation and excellent numerical performance, many other versions of the CG method have been proposed (see [Polak1969, Dai2001, Fletcher1987, Dai1999] and the references therein).
+In general, CG methods use the iterative scheme (2), where the search direction $d_{n}$ is defined by
 $d_{n}=\begin{cases}-\nabla f(x_{0}),&\text{ if }n=0,\\ -\nabla f(x_{n})+\beta_{n-1}d_{n-1},&\text{ if }n\geq 1,\end{cases}$ (4)
-where $\beta_{n}$ is the CG parameter. There are many formulas for $\beta_{n}$, each resulting in different computational behavior [7].
-Based on an extended conjugacy condition, one of the fundamental CG methods was proposed by Dai and Liao [4] with the CG parameter defined as:
+where $\beta_{n}$ is the CG parameter. There are many formulas for $\beta_{n}$, each resulting in different computational behavior [HagerZhang2006].
+One of the most well-known CG methods was proposed by Dai and Liao [Dai2001]. Their approach is based on an extended conjugacy condition, where the CG parameter is defined by
 $\beta_{n}^{DL}=\frac{\langle\nabla f(x_{n+1}),y_{n}\rangle}{\langle d_{n},y_{n}\rangle}-t\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},y_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0},$ (5)
 where $t\in(0,\infty)$, $s_{n}=x_{n+1}-x_{n}$, and $y_{n}=\nabla f(x_{n+1})-\nabla f(x_{n})$.
-Note that when $t=0$, $\beta_{n}^{DL}$ simplifies to the CG parameter proposed by Hestenes and Stiefel [1]:
+Note that when $t=0$, $\beta_{n}^{DL}$ simplifies to the CG parameter proposed by Hestenes and Stiefel [Hestenes1952]:
 $\beta_{n}^{HS}=\frac{\langle\nabla f(x_{n+1}),y_{n}\rangle}{\langle d_{n},y_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0}.$
-The CG parameter proposed by Hager and Zhang [8] is defined as:
-$\beta_{n}^{HZ}=\frac{\langle\nabla f(x_{n+1}),y_{n}\rangle}{\langle d_{n},y_{n}\rangle}-2\frac{\|y_{n}\|^{2}}{\langle d_{n},y_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},y_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0},$
-which can be viewed as an adaptive version of (5) for the specific choice $t=2\frac{\|y_{n}\|^{2}}{\langle d_{n},y_{n}\rangle}$. Similarly, the CG parameter proposed by Dai and Kou [9] is defined as:
+The CG parameter proposed by Hager and Zhang [HagerZhang2005] is defined by
+$\beta_{n}^{HZ}=\frac{\langle\nabla f(x_{n+1}),y_{n}\rangle}{\langle d_{n},y_{n}\rangle}-2\frac{\|y_{n}\|^{2}}{\langle d_{n},y_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},y_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0}.$
+The parameter $\beta_{n}^{HZ}$ can be viewed as an adaptive version of (5) for the specific choice $t=2\frac{\|y_{n}\|^{2}}{\langle d_{n},y_{n}\rangle}$.
+Similarly, the CG parameter proposed by Dai and Kou [DaiKou2013] is defined by
 $\beta_{n}^{DK}(\tau_{n})=\frac{\langle\nabla f(x_{n+1}),y_{n}\rangle}{\langle d_{n},y_{n}\rangle}-\left(\tau_{n}+\frac{\|y_{n}\|^{2}}{\langle s_{n},y_{n}\rangle}-\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}}\right)\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},y_{n}\rangle},$
-where $\tau_{n}$ is a parameter corresponding to the scaling factor in the scaled memoryless Broyden-Fletcher-Goldfarb-Shanno (BFGS) method, which can be regarded as another adaptive form of (5).
-Inspired by Shanno’s matrix perspective on CG algorithms [10], Babaie-Kafaki and Ghanbari [11] observed that the Dai-Liao search directions can be expressed as $d_{n+1}=-P_{n+1}\nabla f(x_{n+1})$ for all $n\in\mathbb{N}_{0}$, where $P_{n+1}=I-\frac{s_{n}y_{n}^{T}}{\langle s_{n},y_{n}\rangle}+t\frac{s_{n}s_{n}^{T}}{\langle s_{n},y_{n}\rangle}$ is a nonsingular matrix for $t\in(0,\infty)$ and $\langle s_{n},y_{n}\rangle\neq 0$; see [12].
-Conducting an eigenvalue analysis on a symmetrized version of $P_{n+1}$, defined by $A_{n+1}=\frac{P_{n+1}+P_{n+1}^{T}}{2}$, we can derive a two-parameter formula for $t$ in the Dai-Liao CG parameter (5), given by
+where $\tau_{n}$ is a parameter that plays the role of a scaling factor in the scaled memoryless Broyden–Fletcher–Goldfarb–Shanno (BFGS) method, which can be viewed as another adaptive variant of
+Motivated by Shanno’s matrix interpretation of CG methods [Shanno1978], Babaie-Kafaki and Ghanbari [KafakiGh2014] showed that the Dai–Liao search directions can be written in the form $d_{n+1}=-P_{n+1}\nabla f(x_{n+1})$ for all $n\in\mathbb{N}_{0}$, where $P_{n+1}=I-\frac{s_{n}y_{n}^{T}}{\langle s_{n},y_{n}\rangle}+t\frac{s_{n}s_{n}^{T}}{\langle s_{n},y_{n}\rangle}$ is a nonsingular matrix for $t\in(0,\infty)$ and $\langle s_{n},y_{n}\rangle\neq 0$; see also [KafakiEJOR2014].
+By performing an eigenvalue analysis on the symmetrized version of $P_{n+1}$, defined by $A_{n+1}=\frac{P_{n+1}+P_{n+1}^{T}}{2}$, we can derive a two-parameter formula for $t$ in the Dai-Liao CG parameter (5), given by
 $t_{n}^{p,q}=p\frac{\|y_{n}\|^{2}}{\langle s_{n},y_{n}\rangle}-q\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}}\text{ for all }n\in\mathbb{N}_{0},$ (6)
 which ensures the descent condition (3) for $p\in\left(\frac{1}{4},\infty\right)$ and $q\in\left(-\infty,\frac{1}{4}\right)$.
-However, in cases where $\frac{\|y_{n}\|^{2}}{\langle s_{n},y_{n}\rangle}$ and $\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}}$ are nearly equal, their subtraction may not be numerically stable [13].
-Moreover, the parameter $t$ in CG parameter $\beta_{n}^{DL}$ must be nonnegative.
+However, in cases where $\frac{\|y_{n}\|^{2}}{\langle s_{n},y_{n}\rangle}$ and $\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}}$ are nearly equal, their subtraction may not be numerically stable [Kafaki2016].
+Moreover, the parameter $t$ in the CG parameter $\beta_{n}^{DL}$ must be nonnegative.
 To address these issues, they restricted $q$ in (6) to be nonpositive and proposed the following CG parameter:
@@
 Similarly, if $(p,q)=(1,0)$, then $\beta_{n}^{DDL}$ is equivalent to $\beta_{n}^{DK}(\tau_{n})$ with the optimal choice of $\tau_{n}=\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}}$.
+For additional variants of the Dial–Liao conjugate gradient method, we refer to [Kafaki2025, Kafaki2025Opt].
 Another significant category of iterative methods used for solving unconstrained optimization problems is the spectral gradient method.
-Initially introduced by Barzilai and Borwein [14], these methods were later advanced by Raydan [15], who developed a spectral gradient approach specifically for large-scale unconstrained optimization challenges.
+Initially introduced by Barzilai and Borwein [Barzilai1988], these methods were later advanced by Raydan [Raydon1997], who developed a spectral gradient approach specifically for large-scale unconstrained optimization challenges.
 The spectral conjugate direction is defined by
 $d_{n}=\begin{cases}-\nabla f(x_{0}),&\text{ if }n=0,\\ -\theta_{n}\nabla f(x_{n})+\beta_{n-1}d_{n-1},&\text{ if }n\geq 1,\end{cases}$
-where $\{\theta_{n}\}\subset(0,\infty)$ is spectral parameter and $\{\beta_{n}\}\subset(0,\infty)$ is conjugate parameter.
-Building on the encouraging numerical results associated with spectral gradient methods [14, 15], Birgin and Martínez [16] proposed the integration of spectral gradient techniques with conjugate gradient methods, culminating in the development of the first spectral conjugate gradient (CG) method.
+where $\{\theta_{n}\}\subset(0,\infty)$ is the spectral parameter and $\{\beta_{n}\}\subset(0,\infty)$ is the conjugate parameter.
+Building on the encouraging numerical results associated with spectral gradient methods [Barzilai1988, Raydon1997], Birgin and Martínez [Birgin2001] proposed the integration of spectral gradient techniques with conjugate gradient methods, culminating in the development of the first spectral conjugate gradient (CG) method.
 The spectral CG methods represent a powerful enhancement of traditional CG methods by incorporating the principles of spectral gradients within the conjugate gradient framework.
-Furthering this line of research, Yu et al.
-[17] modified the spectral Perry’s CG method [18], resulting in an efficient spectral CG variant.
-Capitalizing on the robust convergence properties of quasi-Newton methods, Andrei [19] introduced a scaled BFGS preconditioned CG method in 2007, leveraging quasi-Newton techniques to enhance the effectiveness of conjugate gradient methods.
-This foundational idea has been elaborated upon by Andrei [20, 21, 22], as well as Babaie-Kafaki and Mahdavi-Amiri [23].
-Using the modified secant condition proposed by Li and Fukushima [24], Zhou and Zhang [25] established the following modified secant condition:
-$\nabla^{2}f(x_{n+1})s_{n}=z_{n},$
+Continuing in this direction, Yu and co-authors [Yu2008OMS] adapted Perry’s spectral CG formulation [Perry1978], obtaining a more effective variant of the method.
+Inspired by the good convergence of quasi-Newton techniques, Andrei [Andrei2007] proposed a CG method using a scaled BFGS preconditioner, and this idea was further developed in later works [Andrei2008, Andrei2010EJOR, Babaie2013MMA].
+Li and Fukushima [Li2001] proposed a modified form of the secant condition, and Zhou and Zhang [Zhou2006] later built on their idea to obtain the following modified secant relation:
+$\nabla^{2}f(x_{n+1})s_{n}=z_{n},$ (7)
 where
 $z_{n}=y_{n}+h_{n}\|\nabla f(x_{n})\|^{r}s_{n},\quad h_{n}=\nu+\max\left\{-\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}},0\right\}\|\nabla f(x_{n})\|^{-r}$ (8)
-with $r,\nu\in(0,\infty)$. The modified secant condition (7) plays a crucial role in establishing the global convergence of the MBFGS method for non-convex functions; see [25].
-It is worth noting that the following relation holds independently of the line search:
+with $r,\nu\in(0,\infty)$.
+The modified secant condition (8) is essential to prove the global convergence of the MBFGS algorithm when dealing with non-convex objectives [Zhou2006].
+Note that the following expression holds irrespective of the line search used.
 $\langle z_{n},s_{n}\rangle$ $=\langle y_{n},s_{n}\rangle+\nu\|\nabla f(x_{n})\|^{r}\|s_{n}\|^{2}+\max\left\{-\frac{\langle s_{n},y_{n}\rangle}{\|s_{n}\|^{2}},0\right\}\|s_{n}\|^{2}$
 $\geq\nu\|\nabla f(x_{n})\|^{r}\|s_{n}\|^{2}>0\text{ for all }n\in\mathbb{N}_{0}.$
-Inspired from the appropriate theoretical properties given by (7), Faramarzi and Amini [26] modified the spectral CG method proposed by Jian et al.
-[27] and considered the following conjugate parameter
+Since (8) leads to favorable theoretical properties, Faramarzi and Amini [Faramarzi2019] modified the spectral CG method of Jian et al.
+[Jian2017OMS] and proposed the following choice for the conjugate parameter:
 $\beta_{n}^{ZDK}=\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}-\frac{\|z_{n}\|^{2}}{\langle d_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}\text{ for all }n\in\mathbb{N}_{0}.$
-Based on this conjugate parameter, they construct the following spectral parameter:
+Based on this conjugate parameter, they constructed the following spectral parameter:
 $\theta_{n+1}=\begin{cases}1-\frac{\|z_{n}\|^{2}}{\langle d_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle\nabla f(x_{n+1}),z_{n}\rangle},&\text{ if }1-\frac{\|z_{n}\|^{2}}{\langle d_{n},z_{n}\rangle}\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle\nabla f(x_{n+1}),z_{n}\rangle}\in\left[\frac{1}{4p}+|q|+\eta,\tau\right],\\ 1,&\text{ otherwise},\end{cases}$
-where $\eta$ and $\tau$ are positive constants. To ensure the global convergence of this method, Faramarzi and Amini [26] used the line search satisfying the strong Wolfe conditions:
+where $\eta$ and $\tau$ are positive constants. To ensure global convergence of this method, Faramarzi and Amini [Faramarzi2019] used the line search that satisfied the strong Wolfe conditions
 $f(x_{n}+\alpha_{n}d_{n})\leq f(x_{n})+\delta\alpha_{n}\langle\nabla f(x_{n}),d_{n}\rangle,$ (9)
 $|\langle\nabla f(x_{n}+\alpha_{n}d_{n}),d_{n}\rangle|\leq-\sigma\langle\nabla f(x_{n}),d_{n}\rangle,$ (10)
-where $10<\delta<\sigma<$ 1 and the following assumptions:
+where $0<\delta<\sigma<1$, and the following assumptions:
   1. (A1)
-    The level set
-    $C_{0}=\{x\in\mathbb{R}^{n}:f(x)\leq f(x_{0})\}$
-    is bounded; namely, there exists a constant $M>0$ such that
-    $\|x\|\leq M\text{ for all }x\in C_{0}.$
+    The level set $C_{0}=\{x\in\mathbb{R}^{n}:f(x)\leq f(x_{0})$ is bounded.
   2. (A2)
-    In some neighbourhood $\Omega$ of $C_{0}$, $f$ is continuously differentiable and its gradient is Lipschitz continuous, namely there exists a constant $L>0$ such that
-    $\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|\text{ for all }x,y\in\Omega.$ (10)
-The primary motivation of this study is to introduce a modified spectral conjugate gradient method and to establish its global convergence for general functions under standard assumptions.
-This new method consistently satisfies the sufficient descent property, regardless of the line search used.
-To demonstrate the effectiveness of the proposed approach, we present numerical results that validate its performance and show its superior convergence compared to the modified spectral conjugate gradient method of Faramarzi and Amini [26] and the scaled conjugate gradient method of Mard and Fakhari [28].
-Furthermore, we explore its application in compressed sensing, emphasizing its practical relevance and significant advantages.
+    $f$ is continuously differentiable in some neighbourhood $\Omega$ of $C_{0}$ with $L$-Lipschitz continuous gradient.
+The primary motivation of the present study is to propose a modified spectral conjugate gradient method, guided by an eigenvalue-based analysis and a modified secant condition, and to prove global convergence for general objective functions under typical conditions.This new method always meets the sufficient-descent requirement, independent of the line search strategy applied.
+To demonstrate the effectiveness of our proposed approach, we present numerical results that validate its performance and show its superior convergence compared to the modified spectral conjugate gradient method of Faramarzi and Amini [Faramarzi2019], and the scaled conjugate gradient method of Mard and Fakhari [Mrad2024].
+Furthermore, we explore its application to compressed sensing, emphasizing its practical relevance and significant advantages.
 2 Preliminaries
-In this section, we present some fundamental results from Euclidean space that will be used to establish the convergence of the proposed method.
+In this section we present some fundamental results regarding Euclidean space that will be used to establish the convergence of our proposed method.
 Lemma 1.
-([26]) Let $a,b,\sigma\in\mathbb{R}$. Then
-$(c+\sigma d)^{2}\leq(1+\sigma^{2})(c^{2}+d^{2})\text{ for all }c,d,\sigma\geq 0,$
+([Faramarzi2019]) Let $a,b,\sigma\in\mathbb{R}$. Then
+$(c+\sigma d)^{2}\leq(1+\sigma^{2})(c^{2}+d^{2})\text{ for all }c,d,\sigma\geq 0.$
 Lemma 2.
-([26]) Let $\mathbb{R}^{m}$ be the Euclidean space. Then
+([Faramarzi2019]) Let $\mathbb{R}^{m}$ be a Euclidean space. Then
 $\langle u,v\rangle\leq\frac{1}{2}(\|u\|^{2}+\|v\|^{2})\text{ for all }u,v\in\mathbb{R}^{m}.$
 Lemma 3.
-Let $\mathbb{R}^{m}$ be the Euclidean space and let $\{u_{n}\}$ and $\{v_{n}\}$ be sequences in $\mathbb{R}^{m}$.
-Let $b_{0},\sigma\in(0,\infty)$ and $u_{0}\in\mathbb{R}^{m}$ and let $\{a_{n}\}$ and $\{b_{n}\}$ be sequences in $(0,\infty)$ satisfying the following conditions:
+Let $\mathbb{R}^{m}$ be a Euclidean space, and let $\{u_{n}\}$ and $\{v_{n}\}$ be sequences in $\mathbb{R}^{m}$.
+Let $\sigma\in(0,\infty)$ and let $\{a_{n}\}$ and $\{b_{n}\}$ be two sequences in $(0,\infty)$ satisfying the following conditions:
   1. (C1)
-    $a_{n}>a$ for all $n\in\mathbb{N}$,
+    $a_{n}>a>0$ for all $n\in\mathbb{N}$,
   2. (C2)
-    $\langle v_{n+1},u_{n+1}\rangle<0$ for all $n\in\mathbb{N}$,
+    $\langle v_{n},u_{n}\rangle<0$ for all $n\in\mathbb{N}$,
   3. (C3)
@@
 Then
-$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}+a_{n+1}^{2}\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}\|v_{n+1}\|^{2}}\geq\frac{a}{1+\sigma^{2}}\frac{\|v_{n+1}\|^{2}}{\|u_{n}\|^{2}}.$
+$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}\geq\frac{\|v_{n+1}\|^{4}}{\|u_{n+1}\|^{2}}\left[\frac{a^{2}}{1+\sigma^{2}}-a_{n+1}^{2}\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}\|v_{n+1}\|^{2}}\right].$
 Proof.
-From (11) and condition (C2), we obtain
+Using (11) and condition (C2), we obtain
 $b_{n}^{2}\|u_{n}\|^{2}$ $=\|u_{n+1}\|^{2}+2a_{n+1}\langle v_{n+1},u_{n+1}\rangle+a_{n+1}^{2}\|v_{n+1}\|^{2}$
@@
 which implies that
-$\frac{\|u_{n+1}\|^{2}}{\|u_{n}\|^{2}}\geq b_{n}^{2}-a^{2}\frac{\|v_{n+1}\|^{2}}{\|u_{n}\|^{2}}.$ (12)
-Again, from (11), we have
-$-\langle v_{n+1},u_{n+1}\rangle+b_{n}\langle v_{n+1},u_{n}\rangle=a_{n+1}\|v_{n+1}\|^{2}.$ (13)
-From condition (C3), we have
+$\frac{\|u_{n+1}\|^{2}}{\|u_{n}\|^{2}}\geq b_{n}^{2}-a_{n+1}^{2}\frac{\|v_{n+1}\|^{2}}{\|u_{n}\|^{2}}.$ (12)
+Again, using (11), we have
+$\langle v_{n+1},u_{n+1}\rangle+b_{n}\langle v_{n+1},u_{n}\rangle=a_{n+1}\|v_{n+1}\|^{2}.$ (13)
+It follows from condition (C3) that
 $b_{n}\langle v_{n+1},u_{n}\rangle\leq b_{n}|\langle v_{n+1},u_{n}\rangle|\leq-\sigma b_{n}\langle v_{n},u_{n}\rangle\leq\sigma b_{n}|\langle v_{n},u_{n}\rangle|.$ (14)
-From (13) and (14), we get
-$-\langle v_{n+1},u_{n+1}\rangle+\sigma b_{n}|\langle u_{n},v_{n}\rangle|\geq a_{n+1}\|v_{n+1}\|^{2}.$
+Using (13) and (14), we get
+$\langle v_{n+1},u_{n+1}\rangle+\sigma b_{n}|\langle u_{n},v_{n}\rangle|\geq a_{n+1}\|v_{n+1}\|^{2}.$
 Squaring both sides and applying Lemma 1 with $c=-\langle v_{n+1},u_{n+1}\rangle>0$ (from (C2)) and $d=b_{n}|\langle v_{n},u_{n}\rangle|>0$ gives
 $\langle v_{n+1},u_{n+1}\rangle^{2}+b_{n}^{2}\langle v_{n},u_{n}\rangle^{2}\geq\frac{a_{n+1}^{2}}{1+\sigma^{2}}\|v_{n+1}\|^{4}.$
-Since $a_{n+1}>a$, this yields that
-$\langle v_{n+1},u_{n+1}\rangle^{2}+b_{n}^{2}\langle v_{n},u_{n}\rangle^{2}$ $\geq\frac{a}{1+\sigma^{2}}\|v_{n+1}\|^{4},$
+Since $a_{n+1}>a$, it follows that
+$\langle v_{n+1},u_{n+1}\rangle^{2}+b_{n}^{2}\langle v_{n},u_{n}\rangle^{2}$ $\geq\frac{a^{2}}{1+\sigma^{2}}\|v_{n+1}\|^{4},$
 which implies that
@@
 $=\frac{1}{\|u_{n+1}\|^{2}}\left[\langle v_{n+1},u_{n+1}\rangle^{2}+\frac{\|u_{n+1}\|^{2}}{\|u_{n}\|^{2}}\langle v_{n},u_{n}\rangle^{2}\right]$
-$\geq\frac{1}{\|u_{n+1}\|^{2}}\left[\frac{a}{1+\sigma^{2}}\|v_{n+1}\|^{4}+\left(\frac{\|u_{n+1}\|^{2}}{\|u_{n}\|^{2}}-b_{n}^{2}\right)\langle v_{n},u_{n}\rangle^{2}\right].$ (15)
-From (12) and (15), we conclude that
-$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n+1},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}\geq\frac{\|v_{n+1}\|^{4}}{\|u_{n+1}\|^{2}}\left[\frac{a}{1+\sigma^{2}}-a_{n+1}^{2}\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}\|v_{n+1}\|^{2}}\right],$
-which implies that
-$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}+a_{n+1}^{2}\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}\|v_{n+1}\|^{2}}\geq\frac{a}{1+\sigma^{2}}\frac{\|v_{n+1}\|^{2}}{\|u_{n}\|^{2}}.$
+$\geq\frac{1}{\|u_{n+1}\|^{2}}\left[\frac{a^{2}}{1+\sigma^{2}}\|v_{n+1}\|^{4}+\left(\frac{\|u_{n+1}\|^{2}}{\|u_{n}\|^{2}}-b_{n}^{2}\right)\langle v_{n},u_{n}\rangle^{2}\right].$ (15)
+Using (12) and (15), we conclude that
+$\frac{\langle v_{n+1},u_{n+1}\rangle^{2}}{\|u_{n+1}\|^{2}}+\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}}\geq\frac{\|v_{n+1}\|^{4}}{\|u_{n+1}\|^{2}}\left[\frac{a^{2}}{1+\sigma^{2}}-a_{n+1}^{2}\frac{\langle v_{n},u_{n}\rangle^{2}}{\|u_{n}\|^{2}\|v_{n+1}\|^{2}}\right].$
 Lemma 4.
-([29]) Let $f:\mathbb{R}^{n}\to\mathbb{R}$ such that assumptions (A1) and (A2) hold.
-Consider any iterative method of the form (2), where $d_{n}$ is a descent direction and $\alpha_{n}$ is obtained by the strong Wolfe conditions (8) and (9).
-Then
+([Zoutendijk]) Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be such that assumptions (A1) and (A2) hold.
+For any iterative method of the form (2) with $d_{n}$ being a descent direction and $\alpha_{n}$ determined by the strong Wolfe conditions (9) and (10), we have
 $\sum_{n=0}^{\infty}\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}<\infty.$
-From Lemma 4, we have
+It follows from Lemma 4 that
 $\lim_{n\to\infty}\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}=0.$ (16)
 3 A modified descent Dai-Liao spectral conjugate gradient method and its convergence analysis
-In this section, we introduce a modified spectral conjugate gradient (CG) method, inspired by ideas from quasi-Newton methods.
+In this section we introduce a modified spectral conjugate gradient (CG) descent method, inspired by ideas from quasi-Newton methods.
 Using the theoretical properties provided by (8) and the Dai-Liao conjugate gradient parameter (5), we propose the following modified Dai-Liao conjugate parameter:
@@
 $Q_{n+1}=I-\frac{s_{n}z_{n}^{T}}{\langle s_{n},z_{n}\rangle}+t\frac{s_{n}s_{n}^{T}}{\langle s_{n},y_{n}\rangle}.$
-Following the methodology in [11], we propose the following formula for the parameter $t$ in the modified Dai-Liao conjugate parameter (17):
+Following the methodology in [KafakiGh2014], we propose the following formula for the parameter $t$ in the modified Dai-Liao conjugate parameter (17):
 $t_{n}^{p,q}=p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\text{ for all }n\in\mathbb{N}_{0}$ (18)
@@
 Proposition 5.
-Let $f:\mathbb{R}^{m}\to\mathbb{R}$ be a differentiable function and let $\{d_{n}\}$ be a sequence of spectral CG search directions generated by (20). Then, we have the following:
+Let $f:\mathbb{R}^{m}\to\mathbb{R}$ be a differentiable function and let $\{d_{n}\}$ be a sequence of spectral CG search directions generated by (20). Then the following statements hold:
   • (a)
@@
 (a) Note that $\langle\nabla f(x_{0}),d_{0}\rangle=-\|\nabla f(x_{0})\|^{2}$.
-Then the inequality (21) holds for $n=0$.
-We show that the inequality (21) holds for $n>0$.
+Therefore inequality (21) holds for $n=0$.
+We claim that the inequality (21) holds for $n>0$.
 Let $n\in\mathbb{N}_{0}$.
 Note that $p\in\left(\frac{1}{4},\infty\right)$ and $q\in\left(-\infty,\frac{1}{4}\right)$.
-From (18), (19) and (20), we get
-$\langle\nabla f(x_{n+1}),d_{n+1}\rangle$ $=\langle\nabla f(x_{n+1}),-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}^{MDDL}d_{n}\rangle$
-$=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
-$\quad-\frac{t_{n}^{p,q}\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}$
-$=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
-$\quad-\left(p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\right)\frac{\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}$
-$=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
-$\quad-p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}\frac{\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}$
-$\quad+q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\frac{\langle s_{n},\nabla f(x_{n+1})\rangle\langle d_{n},\nabla f(x_{n+1})\rangle}{\langle d_{n},z_{n}\rangle}.$
+For simplicity, let us denote $\nabla f(x_{n+1})=F_{n+1}$.
+Then, using (18), (19) and (20), we get
+$\langle F_{n+1},d_{n+1}\rangle$
+$=\langle F_{n+1},-\theta_{n+1}F_{n+1}+\beta_{n}^{MDDL}d_{n}\rangle$
+$=-\theta_{n+1}\|F_{n+1}\|^{2}+\frac{\langle F_{n+1},z_{n}\rangle\langle F_{n+1},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}-\frac{t_{n}^{p,q}\langle s_{n},F_{n+1}\rangle\langle d_{n},F_{n+1}\rangle}{\langle d_{n},z_{n}\rangle}$
+$=-\theta_{n+1}\|F_{n+1}\|^{2}+\frac{\langle F_{n+1},z_{n}\rangle\langle F_{n+1},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
+$\quad-\left(p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\right)\frac{\langle s_{n},F_{n+1}\rangle\langle d_{n},F_{n+1}\rangle}{\langle d_{n},z_{n}\rangle}$
+$=-\theta_{n+1}\|F_{n+1}\|^{2}+\frac{\langle F_{n+1},z_{n}\rangle\langle F_{n+1},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}-p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}\frac{\langle s_{n},F_{n+1}\rangle\langle d_{n},F_{n+1}\rangle}{\langle d_{n},z_{n}\rangle}$
+$\quad+q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\frac{\langle s_{n},F_{n+1}\rangle\langle d_{n},F_{n+1}\rangle}{\langle d_{n},z_{n}\rangle}.$
 Since $s_{n}=\alpha_{n}d_{n}$, we have
-$\langle\nabla f(x_{n+1}),d_{n+1}\rangle$ $=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
-$\quad-p\frac{\|z_{n}\|^{2}\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle}+q\frac{\langle d_{n},z_{n}\rangle}{\|d_{n}\|^{2}}\frac{\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle}$
-$=-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle\langle\nabla f(x_{n+1}),d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
-$\quad-p\frac{\|z_{n}\|^{2}\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle^{2}}+q\frac{\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\|d_{n}\|^{2}}.$ (22)
-From Lemma 2 with $u_{n}=\frac{1}{\sqrt{2p}}\nabla f(x_{n+1})$ and $v_{n}=\frac{\sqrt{2p}\langle\nabla f(x_{n+1}),d_{n}\rangle z_{n}}{\langle d_{n},z_{n}\rangle},$ we obtain
-$\frac{\langle\nabla f(x_{n+1}),d_{n}\rangle\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}\leq\frac{1}{4p}\|\nabla f(x_{n+1})\|^{2}+\frac{p\|z_{n}\|^{2}\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\langle d_{n},z_{n}\rangle^{2}}.$ (23)
-On the other hand, by Cauchy-Schwarz inequality, we have
-$q\frac{\langle d_{n},\nabla f(x_{n+1})\rangle^{2}}{\|d_{n}\|^{2}}\leq|q|\frac{\|d_{n}\|^{2}\|\nabla f(x_{n+1})\|^{2}}{\|d_{n}\|^{2}}=|q|\|\nabla f(x_{n+1})\|^{2}.$ (24)
+$\langle F_{n+1},d_{n+1}\rangle$ $=-\theta_{n+1}\|F_{n+1}\|^{2}+\frac{\langle F_{n+1},z_{n}\rangle\langle F_{n+1},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
+$\quad-p\frac{\|z_{n}\|^{2}\langle d_{n},F_{n+1}\rangle^{2}}{\langle d_{n},z_{n}\rangle}+q\frac{\langle d_{n},z_{n}\rangle}{\|d_{n}\|^{2}}\frac{\langle d_{n},F_{n+1}\rangle^{2}}{\langle d_{n},z_{n}\rangle}$
+$=-\theta_{n+1}\|F_{n+1}\|^{2}+\frac{\langle F_{n+1},z_{n}\rangle\langle F_{n+1},d_{n}\rangle}{\langle d_{n},z_{n}\rangle}$
+$\quad-p\frac{\|z_{n}\|^{2}\langle d_{n},F_{n+1}\rangle^{2}}{\langle d_{n},z_{n}\rangle^{2}}+q\frac{\langle d_{n},F_{n+1}\rangle^{2}}{\|d_{n}\|^{2}}.$ (22)
+Using Lemma 2 with $u_{n}=\frac{1}{\sqrt{2p}}F_{n+1}$ and $v_{n}=\frac{\sqrt{2p}\langle F_{n+1},d_{n}\rangle z_{n}}{\langle d_{n},z_{n}\rangle},$ we obtain
+$\frac{\langle F_{n+1},d_{n}\rangle\langle F_{n+1},z_{n}\rangle}{\langle d_{n},z_{n}\rangle}\leq\frac{1}{4p}\|F_{n+1}\|^{2}+\frac{p\|z_{n}\|^{2}\langle d_{n},F_{n+1}\rangle^{2}}{\langle d_{n},z_{n}\rangle^{2}}.$ (23)
+On the other hand, by the Cauchy-Schwarz inequality, we have
+$q\frac{\langle d_{n},F_{n+1}\rangle^{2}}{\|d_{n}\|^{2}}\leq|q|\frac{\|d_{n}\|^{2}\|F_{n+1}\|^{2}}{\|d_{n}\|^{2}}=|q|\|F_{n+1}\|^{2}.$ (24)
 Combining (22), (23) and (24), we conclude that
-$\langle\nabla f(x_{n+1}),d_{n+1}\rangle$ $\leq-\theta_{n+1}\|\nabla f(x_{n+1})\|^{2}+\frac{1}{4p}\|\nabla f(x_{n+1})\|^{2}+|q|\|\nabla f(x_{n+1})\|^{2}$
-$=-\left(\theta_{n+1}-\left(\frac{1}{4p}+|q|\right)\right)\|\nabla f(x_{n+1})\|^{2}.$
+$\langle F_{n+1},d_{n+1}\rangle$ $\leq-\theta_{n+1}\|F_{n+1}\|^{2}+\frac{1}{4p}\|F_{n+1}\|^{2}+|q|\|F_{n+1}\|^{2}$
+$=-\left(\theta_{n+1}-\left(\frac{1}{4p}+|q|\right)\right)\|F_{n+1}\|^{2}.$
 (b) Suppose that $\theta_{n}>\frac{1}{4p}+|q|$ for all $n\in\mathbb{N}$.
-Then, from (21), we get $\langle\nabla f(x_{n}),d_{n}\rangle<0$ for all $n\in\mathbb{N}_{0}$.
-Therefore, $\{d_{n}\}$ is a descent direction.
+Then, employing (21), we get $\langle\nabla f(x_{n}),d_{n}\rangle<0$ for all $n\in\mathbb{N}_{0}$.
+Therefore, $\{d_{n}\}$ is a descent direction as claimed.
 ∎
 Now, we construct the spectral parameter for our conjugate gradient method based on the conjugate parameter $\beta_{n}^{MDDL}$.
-Quasi-Newton methods are a prominent class of iterative optimization techniques that do not require computing the exact Hessian matrix at each iteration.
-The primary motivation behind these methods is to achieve the rapid convergence rate of the classical Newton method while reducing the computational cost.
-Instead of calculating the Hessian matrix directly, quasi-Newton methods construct a sequence of Hessian approximations.
-Inspired by the quasi-Newton direction, the search direction $d_{n+1}$ defined by (20) is designed to mimic the behavior of the quasi-Newton method.
-Consequently, we aim to determine a suitable spectral parameter $\theta_{n+1}$ such that
-$-B_{n+1}^{-1}\nabla f(x_{n+1})\approx-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}^{MDDL}d_{n},$
-where $B_{n+1}$ is an approximation of the Hessian matrix $\nabla^{2}f(x_{n+1})$. Taking inner product with $B_{n+1}s_{n}$ gives that
-$-\langle s_{n},\nabla f(x_{n+1})\rangle\approx-\theta_{n+1}\langle B_{n+1}s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle B_{n+1}s_{n},d_{n}\rangle,$
+Quasi-Newton methods are a prominent class of iterative optimization methods that avoid computing the exact Hessian at every step.
+Their main goal is to retain the fast convergence associated with Newton’s method while keeping the computational burden low.
+Rather than evaluating the true Hessian, these methods update an approximate one at each iteration.
+Motivated by the structure of quasi-Newton directions, the search direction $d_{n}$ in (20) is constructed to emulate this behavior.
+Our objective, therefore, is to choose an appropriate spectral parameter $\theta_{n+1}$ such that
+$B_{n+1}^{-1}\nabla f(x_{n+1})\approx-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}^{MDDL}d_{n},$
+where $B_{n+1}$ is an approximation of the Hessian matrix $\nabla^{2}f(x_{n+1})$. Taking the inner product with $B_{n+1}s_{n}$, we find that
+$\langle s_{n},\nabla f(x_{n+1})\rangle\approx-\theta_{n+1}\langle B_{n+1}s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle B_{n+1}s_{n},d_{n}\rangle,$
 which implies that
-$\theta_{n+1}\approx\frac{1}{\langle B_{n+1}s_{n},\nabla f(x_{n+1})\rangle}(\langle s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle B_{n+1}s_{n},d_{n}\rangle).$
-Since $B_{n+1}$ must satisfies the modified secant condition $B_{n+1}s_{n}=z_{n}$, we have
-${\theta_{n+1}}$ $=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}(\langle s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle z_{n},d_{n}\rangle)$
-$=\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}\left(\langle s_{n},\nabla f(x_{n+1})\rangle+\left(\frac{\langle\nabla f(x_{n+1}),z_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right.\right.$
+$\theta_{n+1}\approx\frac{1}{\langle B_{n+1}s_{n},\nabla f(x_{n+1})\rangle}(\langle s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle B_{n+1}s_{n},d_{n}\rangle).$ (25)
+From (7), $B_{n+1}$ must satisfy the modified secant condition $B_{n+1}s_{n}=z_{n}$. It follows from (25) that
+${\theta_{n+1}}$ $\approx\frac{1}{\langle z_{n},\nabla f(x_{n+1})\rangle}(\langle s_{n},\nabla f(x_{n+1})\rangle+\beta_{n}^{MDDL}\langle z_{n},d_{n}\rangle)$
+[…]
 $\quad\left.\left.-t_{n}^{p,q}\frac{\langle\nabla f(x_{n+1}),s_{n}\rangle}{\langle d_{n},z_{n}\rangle}\right)\langle z_{n},d_{n}\rangle\right)$
@@
 $=1-(t_{n}^{p,q}-1)\frac{\langle s_{n},\nabla f(x_{n+1})\rangle}{\langle z_{n},\nabla f(x_{n+1})\rangle}.$ (27)
-We neglect $\langle s_{n},\nabla f(x_{n+1})\rangle$ in (25), which leads to
+Neglecting $\langle s_{n},\nabla f(x_{n+1})\rangle$ in (26), we arrive at
 ${\theta_{n+1}}=1-\frac{t_{n}^{p,q}\langle s_{n},\nabla f(x_{n+1})\rangle}{\langle z_{n},\nabla f(x_{n+1})\rangle}.$ (28)
 We denote the $\theta_{n+1}$ obtained in equation (27) by $\theta_{n+1}^{R}$ and the $\theta_{n+1}$ obtained in equation (28) by $\theta_{n+1}^{N}$.
-To obtain the sufficient descent condition and the bounded property of the spectral parameters, we propose the following spectral parameters:
+To ensure that the sufficient-descent condition holds and to keep the spectral parameters bounded, we introduce the following choices for the spectral parameters:
 $\theta_{n+1}=\begin{cases}\theta_{n+1}^{R},&\text{ if }\theta_{n+1}^{R}\in\left[\frac{1}{4p}+|q|+\eta,\tau\right],\\ 1,&\text{ otherwise }\end{cases}$ (29)
@@
 $\theta_{n+1}=\begin{cases}\theta_{n+1}^{N},&\text{ if }\theta_{n+1}^{N}\in\left[\frac{1}{4p}+|q|+\eta,\tau\right],\\ 1,&\text{ otherwise },\end{cases}$ (30)
-where $\eta$ and $\tau$ are positive constants. Relation (28) and (29) lead to
+where $\eta$ and $\tau$ are positive constants. Relations (29) and (30) lead to
 $\frac{1}{4p}+|q|+\eta\leq\theta_{n+1}\leq\tau\textrm{ for all }n\in\mathbb{N}_{0}.$ (31)
 Now, we propose the following modified descent Dai-Liao spectral conjugate gradient method (MDDLSCG):
-Input: Let $x_{0}\in\mathbb{R}^{n}$ and let $\epsilon,\delta,\sigma,\eta,\tau,r,\nu\in(0,\infty)$ such that
-     $10<\delta<\sigma<$ 1.
+Input: Let $x_{0}\in\mathbb{R}^{n}$ and let $\epsilon,\delta,\sigma,\eta,\tau,r,\nu\in(0,\infty)$ be such that
+     $0<\delta<\sigma<1$.
 Step 0: Set $d_{0}=-\nabla f(x_{0})$.
-Step 1: If $\|\nabla f(x_{n})\|_{\infty}<\epsilon$, stop.
+Step 1: If $\|\nabla f(x_{n})\|_{\infty}<\varepsilon$, stop, where $\|u\|_{\infty}=\max_{1\leq i\leq m}|u_{i}|$ for
+    $u=(u_{1},\ldots,u_{m})\in\mathbb{R}^{m}$.
 Step 2: Determine the step length $\alpha_{n}$ satisfying the strong Wolfe
@@
 Step 4: Compute $d_{n+1}$ by
-$d_{n+1}=-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}d_{n}\text{ for all }n\in\mathbb{N}_{0},$
-      where CG parameter $\beta_{n}^{MDDL}$ is given by (19) and spectral
-    parameter is given by (28) or (29).
+$d_{n+1}=-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}^{MDDL}d_{n}\text{ for all }n\in\mathbb{N}_{0},$
+    where the CG parameter $\beta_{n}^{MDDL}$ is given by (19) and
+      the spectral parameter $\theta_{n}$ is given by (29) or (30).
 Step 5: Set $n=n+1$ and return to Step 1.
@@
 (i) Using $\theta_{n}\geq\frac{1}{4p}+|q|+\eta$ and (21), we conclude that
-$\langle\nabla f(x_{n}),d_{n}\rangle\leq-\eta\|\nabla f(x_{n})\|^{2}\text{ for all }n\in\mathbb{N}.$ (31)
-(ii) If $\|\nabla f(x_{n})\|\neq 0$, then, from (9) and (31), we see that
+$\langle\nabla f(x_{n}),d_{n}\rangle\leq-\eta\|\nabla f(x_{n})\|^{2}\text{ for all }n\in\mathbb{N}_{0}.$ (32)
+(ii) If $\|\nabla f(x_{n})\|\neq 0$, then using (10) and (32), we see that
 $\langle d_{n},y_{n}\rangle\geq-(1-\sigma)\langle\nabla f(x_{n}),d_{n}\rangle\geq(1-\sigma)\eta\|\nabla f(x_{n})\|^{2}>0,$ (33)
-which, from (7), implies that $h_{n}=\nu$ and hence
+which, by (8), implies that $h_{n}=\nu$ and hence
 $z_{n}=y_{n}+\nu\|\nabla f(x_{n})\|^{r}s_{n}.$ (34)
@@
   1. (i)
-    From (8) and (31), we can see that $\{f(x_{n})\}$ is decreasing. Thus, the sequence $\{x_{n}\}$ generated by Algorithm 1 is contained in $C_{0}$.
+    Employing (9) and (32), we can see that $\{f(x_{n})\}$ is decreasing. Thus, the sequence $\{x_{n}\}$ generated by Algorithm 1 is contained in $C_{0}$.
   2. (ii)
-    From Assumption (A2), we have
+    From Assumption (A1), there exists $M>0$ such that $\|x\|\leq M$ for all $x\in C_{0}$. Then using Assumption (A2), we have
     $\|\nabla f(x)\|$ $\leq\|\nabla f(x)-\nabla f(x_{0})\|+\|\nabla f(x_{0})\|$
     $\leq L\|x-x_{0}\|+\|\nabla f(x_{0})\|$
-    $\leq 2ML+\|\nabla f(x_{0})\|:=\gamma\text{ for all }x\in C_{0}.$ (34)
+    $\leq 2ML+\|\nabla f(x_{0})\|=\gamma\text{ for all }x\in C_{0}.$ (35)
 In addition, suppose that $\nabla f(x_{n})\neq 0$ for all $n\in\mathbb{N}_{0}$; otherwise, a stationary point will be found.
@@
 Proof.
-Since $\liminf_{n\to\infty}\|\nabla f(x_{n})\|\neq 0$. Then, there exists $\epsilon>0$ such that $\|\nabla f(x_{n})\|\geq\epsilon$ for all $n\in\mathbb{N}_{0}$. This together with (16) implies that
+Since $\liminf_{n\to\infty}\|\nabla f(x_{n})\|\neq 0$, there exists $\epsilon>0$ such that $\|\nabla f(x_{n})\|\geq\epsilon$ for all $n\in\mathbb{N}_{0}$. When combined with (16), this implies that
 $\lim_{n\to\infty}\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}\|\nabla f(x_{n+1})\|^{2}}=0.$ (36)
 Let $u_{n}=d_{n}$, $v_{n}=\nabla f(x_{n})$, $a_{n}=\theta_{n}$ and $b_{n}=\beta_{n}$ in Lemma 3.
-Since $\theta_{n}>\frac{1}{4p}+|q|$, it follows from Remark 6 that $\langle\nabla f(x_{n}),d_{n}\rangle<0$ for all $n\in\mathbb{N}_{0}$.
-Also from strong Wolfe condition (9), we have $|\langle\nabla f(x_{n+1}),d_{n}\rangle|\leq-\sigma\langle\nabla f(x_{n}),d_{n}\rangle$.
+Since $\theta_{n}>\frac{1}{4p}+|q|>0$, it follows from Remark 6 that $\langle\nabla f(x_{n}),d_{n}\rangle<0$ for all $n\in\mathbb{N}_{0}$.
+Also from strong Wolfe condition (10), we have $|\langle\nabla f(x_{n+1}),d_{n}\rangle|\leq-\sigma\langle\nabla f(x_{n}),d_{n}\rangle$ for all $n\in\mathbb{N}_{0}$.
 Then, clearly all the conditions (C1), (C2) and (C3) of Lemma 3 are satisfied.
 Thus, we have
-$\frac{\langle\nabla f(x_{n+1}),d_{n+1}\rangle^{2}}{\|d_{n+1}\|^{2}}+\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}\geq b_{1}\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}-\theta_{n+1}^{2}\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}\|\nabla f(x_{n+1})\|^{2}},$
-where $b_{1}=\left(\frac{1}{4p}+|q|\right)\frac{1}{1+\sigma^{2}}$. It follows from (35) that for all sufficiently large $n$, we have
-$\frac{\langle\nabla f(x_{n+1}),d_{n+1}\rangle^{2}}{\|d_{n+1}\|^{2}}+\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}\geq b_{1}\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}.$ (36)
-From (36) and Lemma 4, we get that $\sum_{n=1}^{\infty}\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}<\infty.$
+$\frac{\langle\nabla f(x_{n+1}),d_{n+1}\rangle^{2}}{\|d_{n+1}\|^{2}}+\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}\geq\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}\left[B-\theta_{n+1}^{2}\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}\|\nabla f(x_{n+1})\|^{2}}\right],$
+where $B=\left(\frac{1}{4p}+|q|\right)^{2}\frac{1}{1+\sigma^{2}}$. It follows from (36) that for all sufficiently large $n$, we have
+$\frac{\langle\nabla f(x_{n+1}),d_{n+1}\rangle^{2}}{\|d_{n+1}\|^{2}}+\frac{\langle\nabla f(x_{n}),d_{n}\rangle^{2}}{\|d_{n}\|^{2}}\geq B\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}.$ (37)
+Using (37) and Lemma 4, we get that $\sum_{n=1}^{\infty}\frac{\|\nabla f(x_{n+1})\|^{4}}{\|d_{n+1}\|^{2}}<\infty.$
 Theorem 9.
-Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a function and let $\{x_{n}\}$ be the sequence generated by Algorithm 1 such that Assumptions (A1) and (A2) hold.
+Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be a real function and let $\{x_{n}\}$ be the sequence generated by Algorithm 1 such that Assumptions (A1) and (A2) hold.
 Then $\liminf_{n\to\infty}\|\nabla f(x_{n})\|=0.$
 Proof.
-Suppose the claim of the theorem is not true, i.e., there exists a positive constant $\xi$ such that
+Suppose, for the sake of contradiction, that the assertion of the theorem does not hold. Then there exists a positive constant $\xi$ such that
 $\|\nabla f(x_{n})\|>\xi\text{ for all }n\in\mathbb{N}_{0}.$ (38)
-From (32), (33) and (37), we obtain
+Employing (33), (34) and (38), we obtain
 $\langle d_{n},z_{n}\rangle$ $=\langle d_{n},y_{n}\rangle+\nu\|\nabla f(x_{n})\|^{r}\alpha_{n}\|d_{n}\|^{2}$
 $\geq\nu\|\nabla f(x_{n})\|^{r}\alpha_{n}\|d_{n}\|^{2}\geq\nu\xi^{r}\alpha_{n}\|d_{n}\|^{2}.$ (39)
-From (33), (10) and ((ii)), we have
+Using Assumption (A2), (34) and ((ii)), we have
 $\|z_{n}\|\leq\|y_{n}\|+\nu\|\nabla f(x_{n})\|^{r}\|s_{n}\|\leq L\|s_{n}\|+\nu\gamma^{r}\|s_{n}\|=(L+\nu\gamma^{r})\|s_{n}\|.$ (40)
-From the Lipschitz continuity of the $\nabla f$ and (39), we obtain
-$t_{n}^{p,q}$ $\leq\left|p\frac{\|z_{n}\|^{2}}{\langle s_{n},z_{n}\rangle}-q\frac{\langle s_{n},z_{n}\rangle}{\|s_{n}\|^{2}}\right|$
-$\leq p\frac{(L+\nu\gamma^{r})^{2}\|s_{n}\|^{2}}{\mu\|s_{n}\|^{2}}+|q|\frac{(L+\nu\gamma^{r})\|s_{n}\|^{2}}{\|s_{n}\|^{2}}=p\frac{L^{2}}{\mu}+L|q|.$
-Thus, the relations (19), (10), ((ii)), (38), (39), along with Cauchy-Schwarz inequality, conclude
+Thus, relations (19), ((ii)), (39), (40), along with the Cauchy-Schwarz inequality, yield
 $\|\beta_{n}^{MDDL}\|$
@@
 $\leq\frac{\gamma(L+\nu\gamma^{r})}{\nu\xi^{r}\|d_{n}\|}+\frac{p\gamma(L+\nu\gamma^{r})^{2}}{\nu^{2}\xi^{2r}\|d_{n}\|}+|q|\frac{\gamma}{\|d_{n}\|}.$
-Note that $\theta_{n+1}\leq\tau$ by (30). Hence, from (20), (40), we get
+Note that $\theta_{n+1}\leq\tau$ by (31). Hence, using (20) and (41), we arrive at
 $\|d_{n+1}\|$ $=\|-\theta_{n+1}\nabla f(x_{n+1})+\beta_{n}^{MDDL}d_{n}\|$
@@
 $\leq\tau\gamma+\frac{\gamma(L+\nu\gamma^{r})}{\nu\xi^{r}}+\frac{p\gamma(L+\nu\gamma^{r})^{2}}{\nu^{2}\xi^{2r}}+|q|\gamma.$ (42)
-From (41), we observe that
+Using (42), we observe that
 $\frac{1}{\|d_{n+1}\|^{2}}\geq\frac{1}{E^{2}}\textrm{ for all }n\in\mathbb{N}_{0},$
@@
 $\sum_{n=0}^{\infty}\frac{1}{\|d_{n}\|^{2}}=\infty.$ (43)
-From (37) and Lemma 8, we have
+Employing (38) and Lemma 8, we find that
 $\sum_{n=0}^{\infty}\frac{1}{\|d_{n}\|^{2}}=\sum_{n=0}^{\infty}\frac{\|\nabla f(x_{n})\|^{4}}{\|d_{n}\|^{2}\|\nabla f(x_{n})\|^{4}}\leq\frac{1}{\xi^{4}}\sum_{n=0}^{\infty}\frac{\|\nabla f(x_{n})\|^{4}}{\|d_{n}\|^{2}}<\infty,$
-which contradicts (42). Therefore, $\liminf_{n\to\infty}\|\nabla f(x_{n})\|=0$. ∎
+which contradicts (43). Therefore, $\liminf_{n\to\infty}\|\nabla f(x_{n})\|=0$, as asserted. ∎
 4 Numerical Experiments
-In this section, we discuss the numerical performance of our proposed method (MDDLSCG), comparing it with Algorithm 3.1 in [26] (denoted by MSCG) and Scaled conjugate gradient method (denoted by ScCG) in [28].
+In this section we present the numerical performance of our proposed method (MDDLSCG), comparing it with Algorithm 3.1 in [Faramarzi2019] (denoted by MSCG) and the scaled conjugate gradient method (denoted by ScCG) in [Mrad2024].
 All numerical experiments were conducted in MATLAB R2024b on an HP workstation with Intel Core i7-13700 processor, 32 GB RAM and Windows 11.
-For the numerical experiment, we consider two test problems: (i) finding the minima of the Beale function, and (ii) finding the minima of the Almost Perturbed Quadratic (APQ) function:
+For the numerical experiments, we consider two test problems: (i) finding the minima of the Beale function, and (ii) finding the minima of the Almost Perturbed Quadratic (APQ) function.
 Beale Test Function: The Beale function is multimodal with sharp peaks at the corners of the input domain, which is defined from the general Beale function as follows:
@@
 The global minimum is located at $(x,y)=(3,0.5)$, where $f(x,y)=0$.
-During the numerical experiments, we set the parameters $\delta=0.01$, $\sigma\in\{0.1,0.2,0.4,0.6,0.8,0.9\}$ for all the algorithms, $\eta=0.001$, $\tau=10$, $r=1$ and $\nu=0.001$ for algorithms MSCG and MDDLSCG and $p=0.4$ and $q=0.2$ for the algorithm MDDLSCG.
+(a) Surface and Contour plot of the Beale functions
+For the numerical tests, we choose $\delta=0.01$, $\sigma\in\{0.1,0.2,0.4,0.6,0.8,0.9\}$ for all the algorithms, $\eta=0.001$, $\tau=10$, $r=1$ and $\nu=0.001$ for algorithms MSCG and MDDLSCG and $p=0.4$ and $q=0.2$ for the algorithm MDDLSCG.
 We take the initial point $x_{0}=(1,0.8)$.
-During the implementation, the iterations are terminated, when it meets the stopping criterion $E_{n}=\|\nabla f(x_{n})\|_{\infty}<10^{-15}$.
-The numerical results of all the algorithms MDDLSCG, MSCG and ScCG for the Beale function are summarized in Table 1, while their performance is illustrated in Figure LABEL:bealefig1 and LABEL:bealefig2.
-In Figures LABEL:bealefig1 (a), (c), (e), the iteration paths are visualized on the contour plot of the Beale function, showing how each method converges to the minimizer for $\sigma=0.1,0.2,0.6$, respectively.
-Figures LABEL:bealefig1 (b), (d), (f) depict the convergence rate by plotting the norm of the gradient, $\|\nabla f(x_{n})\|_{\infty}$ versus the number of iterations on a logarithmic scale for $\sigma=0.1,0.2,0.6$, respectively.
-The graph between different values of sigma and number of iterations is given in Figure LABEL:bealefig2 (a) and the graph between different values of sigma and execution time is in Figure LABEL:bealefig2 (b).
-From the data presented, it is evident that our proposed method, MDDLSCG, outperforms both MSCG and ScCG in all tested scenarios.
-As shown in Table 1, Figure LABEL:bealefig1 and LABEL:bealefig2 (b), MDDLSCG reaches the stopping criteria in fewer iterations and requires less execution time.
-Additionally, From Figure LABEL:bealefig2 (a), we observe that these methods are significantly sensitive to variations in the parameter $\sigma$, which affects their performance.
+The iterations are terminated once they satisfy the stopping criterion $E_{n}=\|\nabla f(x_{n})\|_{\infty}<10^{-15}$.
+The numerical results of all the algorithms MDDLSCG, MSCG and ScCG for the Beale function are summarized in Table 1, while their performance is illustrated in Figure 8(a) and 8(b).
+In Figures 8(a) (a), (c), (e), the iteration paths are visualized on the contour plot of the Beale function, showing how each method converges to the minimizer for $\sigma=0.1,0.2,0.6$, respectively.
+Figures 8(a) (b), (d), (f) depict the convergence rate by plotting the norm of the gradient, $\|\nabla f(x_{n})\|_{\infty}$ versus the number of iterations on a logarithmic scale for $\sigma=0.1,0.2,0.6$, respectively.
+The comparison graph between different values of sigma and the number of iterations is depicted in Figure 8(b) (a) and the comparison graph between different values of sigma and execution time is in Figure 8(b) (b).
+In view of the data presented, it is evident that our proposed method, MDDLSCG, outperforms both MSCG and ScCG in all tested scenarios.
+As shown in Table 1, Figure 8(a) and 8(b) (b), MDDLSCG reaches the stopping criteria in fewer iterations and requires less execution time.
+Additionally, using Figure 8(b) (a), we observe that these methods are significantly sensitive to variations in the parameter $\sigma$, which affects their performance.
 $\sigma$ MDDLSCG MSCG ScCG itr Tcpu $E_{n}$ itr Tcpu $E_{n}$ itr Tcpu $E_{n}$ 0.1 37 1.86e-02 9.49e-15 63 3.14e-02 2.63e-16 200 5.40e-02 5.27e-16 0.2 36 1.75e-02 3.16e-15 63 2.76e-02 9.07e-15 136 3.99e-02 4.51e-15 0.4 53 2.44e-02 9.40e-15 73 2.65e-02 8.75e-15 126 4.73e-02 2.87e-15 0.6 40 2.36e-02 7.91e-16 72 2.78e-02 5.74e-15 95 4.94e-02 9.32e-15 0.8 67 3.24e-02 5.32e-15 111 2.86e-02 4.48e-15 201 4.92e-02 6.91e-15 0.9 64 3.28e-02 8.16e-15 101 3.48e-02 8.74e-15 184 4.37e-02 1.08e-15
 Table 1: Comparison table for MDDLSCG, MSCG and ScCG for Beale function
+(a) Comparison of the iterations and execution time for different $\sigma$ for MDDLSCG, MSCG, ScCG for the Beale function
+(b) Comparison of the iterations and execution time for different $\sigma$ for MDDLSCG, MSCG, ScCG for the Beale function
 Almost Perturbed Quadratic function: Consider the Almost Perturbed Quadratic (APQ) function $f:\mathbb{R}^{m}\to\mathbb{R}$ defined by
-$f(\mathbf{x})=\frac{1}{2}\mathbf{x}^{T}A\mathbf{x}+\mathbf{b}^{T}\mathbf{x}$
-where $A\in\mathbb{R}^{m\times m}$ is a symmetric positive definite matrix ans $b\in\mathbb{R}^{m}$ is a vector of perturbation.
+$f(x)=\frac{1}{2}x^{T}Ax+b^{T}x\text{ for all }x\in\mathbb{R}^{m},$
+where $A\in\mathbb{R}^{m\times m}$ is a symmetric positive definite matrix ans $b\in\mathbb{R}^{m}$ is a vector of perturbations.
 This function has a global minimum at $x=-A^{-1}b$.
@@
 $\nabla f(x)=Ax+b\text{ for all }x\in\mathbb{R}^{m}.$
-For the numerical experiments, we have considered three cases, each characterized by dimensions of1000100010001000, $100$ and $25$.
-To the cases with dimensions $1000$ and $100$, the matrix $A$ and vector $b$ are randomly generated using MATLAB’s built-in function rand and the initial point is randomly generated using MATLAB’s rand function.
-For the case with dimension $25$, we take the initial point
+For the numerical experiments, we have considered three cases, each characterized by dimensions of $1000$, 100100 and $25$.
+In the cases with dimensions $1000$ and $100$, the matrix $A$ and vector $b$ are randomly generated using MATLAB’s built-in function rand and the initial point is randomly generated using MATLAB’s rand function.
+In the case of dimension $25$, we take the initial point
 $x_{0}$ $=(x_{0,1},\ldots,x_{0,i},\ldots,x_{0,25}),\text{ where }x_{0,i}=1\text{ for all }i=1,2,\ldots,25\text{ and}$
 $a$ $=(2,4,3,6,4,12,8,10,5,8,6,11,4,8,3,6,2,12,6,8,4,10,5,2,6),$
-$b$ =(4,52,18,60,24,72,72,50,55,80,78,22,12=(-4,-52,-18,-60,-24,-72,-72,-50,-55,-80,-78,-22,-12,= ( - 4 , - 52 , - 18 , - 60 , - 24 , - 72 , - 72 , - 50 , - 55 , - 80 , - 78 , - 22 , - 12 ,
-80,27,42,24,120,30,120,24,100,60,20,66)$\quad$-80,-27,-42,-24,-120,-30,-120,-24,-100,-60,-20,-66),- 80 , - 27 , - 42 , - 24 , - 120 , - 30 , - 120 , - 24 , - 100 , - 60 , - 20 , - 66 ) ,
+$b$ $=(-4,-52,-18,-60,-24,-72,-72,-50,-55,-80,-78,-22,-12,$
+$\quad-80,-27,-42,-24,-120,-30,-120,-24,-100,-60,-20,-66),$
 $A$ $=\text{diag}(a).$
-For all the cases, the parameters are set as follows: $\delta=0.01$ and $\sigma=0.1$ for all algorithms, while for MSCG and MDDLSCG, we use $\eta=0.001$, $\tau=10$, $r=1$, and $\nu=0.001$.
+In all cases, the parameters are set as follows: $\delta=0.01$ and $\sigma=0.1$ for all algorithms, while for MSCG and MDDLSCG, we use $\eta=0.001$, $\tau=10$, $r=1$, and $\nu=0.001$.
 Additionally, for the MDDLSCG algorithm, we set $p=0.4$ and $q=0.2$.
-During the implementation, the iterations are terminated when the stopping criterion $E_{n}=\|\nabla f(x_{n})\|_{\infty}<10^{-6}$ is satisfied, where $x_{n}=(x_{n,1},\ldots,x_{n,25})$.
+The iterations are terminated once they satisfy the stopping criterion $E_{n}=\|\nabla f(x_{n})\|_{\infty}<10^{-6}$ is satisfied, where $x_{n}=(x_{n,1},\ldots,x_{n,25})$.
 The performance of the algorithms MDDLSCG, MSCG, and ScCG for the APQ function is summarized in Table 2 for all three cases.
-To provide a clearer visualization of the experiment, we have plotted the graph of the number of iterations versus $E_{n}$ in Figure LABEL:APQgraph.
-From Table 2 and Figure LABEL:APQgraph, it is evident that our proposed method, MDDLSCG, requires fewer iterations and less computational time to satisfy the desired stopping criterion across all cases.
-Figure LABEL:APQtrack illustrates how the algorithms MDDLSCG, MSCG, and ScCG approximate the desired solution $x^{*}=-A^{-1}b$ at the 5th and 10th iterations for the case where the problem dimension is $25$.
-Furthermore, Figure LABEL:APQprogress depicts the convergence behavior of the algorithms as the number of iterations increases up to the 50th iteration for the same dimension.
+To provide a clearer visualization of the experiment, we have plotted the graph of the number of iterations versus $E_{n}$ in Figure 11(a).
+From Table 2 and Figure 11(a), it is evident that our proposed method, MDDLSCG, requires fewer iterations and less computational time to satisfy the desired stopping criterion across all cases.
+Figure 11(b) illustrates how the algorithms MDDLSCG, MSCG, and ScCG approximate the desired solution $x^{*}=-A^{-1}b$ at the 5th and 10th iterations in the case where the problem dimension is $25$.
+Furthermore, Figure 11(c) depicts the convergence behavior of the algorithms as the number of iterations increases up to the 50th iteration for the same dimension.
 It is observed that MDDLSCG demonstrates superior convergence characteristics compared to MSCG and ScCG.
@@
 ScCG 171 4.053286e-02 8.747078e-07
-Table 2: Comparison table of MDDLSCG, MSCG and ScCG for APQ function
-5 Application in compressed sensing
-Compressed sensing is a well-known signal processing strategy for efficiently acquiring and reconstructing signals with a sparse structure, allowing for memoryless and compact storage of the signals [30].
-The importance of compressed sensing in real-world applications, such as machine learning, compressive imaging, radar, wireless sensor networks, medical imaging, astrophysical signals, and video coding, has been highlighted in [31].
-The compressed sensing problem involves finding a sparse solution to a highly underdetermined system, represented as $Ax=b$, where $A\in\mathbb{R}^{m\times n}$ (with $m\ll n$) and $b\in\mathbb{R}^{m}$ [30].
+Table 2: Comparison table of MDDLSCG, MSCG and ScCG for the APQ function
+(a) Convergence behavior of $\|\nabla f(x_{n})\|_{\infty}$ for MDDLSCG, MSCG, ScCG for the APQ function
+(b) Tracking behaviour of MDDLSCG, MSCG, ScCG for the APQ function at various iterations
+(c) Graphical progression of MDDLSCG, MSCG, ScCG upto 50th iterations for the APQ function
+5 Application to compressed sensing
+Compressed sensing is a widely recognized signal processing approach that enables the efficient acquisition and reconstruction of signals with a sparse structure, allowing these signals to be stored in a compact and memoryless form [Bruckstein2009].
+Its practical significance has been demonstrated in numerous fields, including machine learning, compressive imaging, radar systems, wireless sensor networks, medical and astronomical signal processing, and video compression, as discussed in [Aminifard2022].
+The compressed sensing problem involves finding a sparse solution to a highly underdetermined system, represented as $Ax=b$, where $A\in\mathbb{R}^{m\times n}$ (with $m\ll n$) and $b\in\mathbb{R}^{m}$ [Bruckstein2009].
 A common approach is to solve the following optimization problem:
 $f(x)=\frac{1}{2}\|Ax-b\|^{2}+\mu\phi(x),$
-where $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is a penalty function, and $\mu>0$ is a parameter that controls the trade-off between sparsity and reconstruction fidelity.
-A popular choice for $\phi$ is $\phi(x)=\|x\|_{1}$, leading to the well-known Basis Pursuit Denoising (BPD) problem, which has been extensively studied in the literature; see [32] and the references therein.
-However, solving BPD is challenging due to the non-smoothness of the $\ell_{1}$-penalty term.
-To address this issue, Zhu et al.
-[33] proposed a relaxation of BPD using Nesterov’s smoothing technique [34], resulting in the following formulation:
+where $\phi:\mathbb{R}^{n}\to\mathbb{R}$ is a penalty function, and $\mu>0$ is a parameter that regulates the trade-off between sparsity and reconstruction fidelity.
+A popular choice for $\phi$ is $\phi(x)=\|x\|_{1}$, leading to the well-known Basis Pursuit Denoising (BPD) problem, which has been extensively studied in the literature; see [Esmaeili2019] and references therein.
+However, solving BPD is difficult because the $\ell_{1}$-penalty term is non-smooth.
+To overcome this difficulty, Zhu et al.
+[Zhu2013] introduced a relaxed form of BPD based on Nesterov’s smoothing technique [Nesterov2005Siam], leading to the following formulation:
 $f_{\lambda}(x)=\sum_{i=1}^{n}\psi_{\lambda}(|x_{i}|),$
-where $x_{i}$ represents the $i$-th component of the vector $x$, and $\psi_{\lambda}$ is defined as:
+where $x_{i}$ represents the $i$-th component of the vector $x$, and $\psi_{\lambda}$ is defined by
 $\psi_{\lambda}(|x_{i}|)=\max_{\gamma\in[0,1]}\left\{\gamma|x_{i}|-\frac{1}{2}\lambda\gamma^{2}\right\},$
 which simplifies to:
-$\psi_{\lambda}(|x_{i}|)=\begin{cases}\frac{|x_{i}|^{2}}{2\lambda},&\text{if }|x_{i}|<\lambda,\\ |x_{i}|-\frac{\lambda}{2},&\text{if }|x_{i}|\geq\lambda.\end{cases}$
-Here, $f_{\lambda}$ is known as the Huber function [35], and its gradient $\nabla f_{\lambda}$ has a Lipschitz constant of $\frac{1}{\lambda}$ [36], with the explicit form given by:
-$(\nabla f_{\lambda})_{i}=\begin{cases}\frac{x_{i}}{\lambda},&\text{if }|x_{i}|<\lambda,\\ \text{sgn}(x_{i}),&\text{otherwise},\end{cases}$
-where $\lambda>0$ is a constant and $\text{sgn}(\cdot)$ denotes the signum function.
-Therefore, instead of solving the original problem $f(x)=\frac{1}{2}\|Ax-b\|^{2}+\mu\|x\|_{1}$, we solve the smoothed version:
+$\psi_{\lambda}(|x_{i}|)=\begin{cases}\frac{|x_{i}|^{2}}{2\lambda},&\text{if }|x_{i}|<\lambda,\\ |x_{i}|-\frac{\lambda}{2},&\text{if }|x_{i}|\geq\lambda,\end{cases}$
+where $\lambda>0$.
+Function $f_{\lambda}$ is known as the Huber function [Rangarajan], and its gradient $\nabla f_{\lambda}$ is $\frac{1}{\lambda}$ -Lipschitz continuous [Becker2011].
+The explicit form of the Huber function is given by
+[…]
+where $\text{sgn}(\cdot)$ denotes the signum function.
+Therefore, instead of solving the original problem for $f(x)=\frac{1}{2}\|Ax-b\|^{2}+\mu\|x\|_{1}$, we solve the smoothed version:
 $\min_{x\in\mathbb{R}^{n}}f(x)=\frac{1}{2}\|Ax-b\|^{2}+\mu f_{\lambda}(x).$
@@
 $\mu=\max\{2^{-7},0.001\|A^{T}b\|_{\infty}\}$, $\lambda=\min\{0.001,0.048\|A^{T}b\|_{\infty}\}$, $\sigma=0.1$, $\delta=0.01$, $\eta=0.001$, $\tau=10$, $r=1$ and $\nu=0.001$.
-[…]
+For MDDLSCG, we choose $p=0.4$ and $q=0.2$.
+The Gaussian matrix $A$ is generated with the MATLAB commands randn(m, n) and the noise vector $w$ is generated with $0.1\times\texttt{randn(m, 1)}$.
+The measurement vector $b$ is computed by $b=Ax+w.$ The regularization parameter is chosen as $\mu=0.001\|A^{T}b\|_{\infty}$.
+Each methods begin with the initial point $x_{0}=A^{T}b$.
+Let $x$ represent the true signal and $x^{*}$ denote the recovered one.
+The termination criterion for the experiment is the mean-squared error(MSE):
 $\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(x_{i}-x_{i}^{*})^{2}\leq 10^{-5}.$
-To evaluate the restoration performance qualitatively, we report the relative error (RelErr) of the recovered signal, which is defined as follows:
+To assess the reconstruction quality, we also compute the relative error (RelErr) of the restored signal, defined by
 $\text{RelErr}=\frac{\|x^{*}-x\|}{\|x\|}.$
-Let $k$ denote the number of nonzero elements in the original signal. For a fair comparison, we consider two different settings: $(m,n,k)=(128,512,16)$, $(256,1024,32)$.
-The detailed numerical results are presented in Table 3, where we report the average number of iterations (Itr), average CPU time (Tcpu), mean squared error (MSE), and relative error (RelErr) for each setting.
-The graph depicting the relationship between relative error and the number of iterations is shown in Figures LABEL:stpp512(a) and LABEL:stopp(a), while the graph illustrating MSE versus the number of iterations is presented in Figures LABEL:stpp512(b) and LABEL:stopp(b).
-Additionally, the recovery of the signal is visualized in Figures 11 and 13.
-From Table 3, Figures LABEL:stpp512, and LABEL:stopp, it is clear that the MDDLSCG algorithm surpasses the MSCG and ScCG algorithms in performance across all cases.
+Let $k$ represent the number of nonzero components in the original signal. To ensure a fair comparison, we test two cases: $(m,n,k)=(128,512,16)$, $(256,1024,32)$.
+The numerical results are summarized in Table 3, where we list, for each case, the average iteration count (Itr), the average CPU time (Tcpu), the mean squared error (MSE), and the relative error (RelErr).
+The graph depicting the relationship between relative error and the number of iterations is shown in Figures 11(d)(a) and 13(a)(a), while the graph illustrating MSE versus the number of iterations is presented in Figures 11(d)(b) and 13(a)(b).
+In addition, the recovery of the signal is visualized in Figures 11 and 13.
+From Table 3, Figures 11(d) and 13(a), it is clear that the MDDLSCG algorithm surpasses the MSCG and ScCG algorithms in performance across all cases.
+(d) Convergence behavior and relative error plots, demonstrating the performance of the methods MDDLSCG, MSCG and ScCG for $(m,n,k)=(128,512,16)$
 Figure 11: Signal recovery for data $(m,n,k)=(128,512,16)$
+(a) Convergence behavior and relative error plots, demonstrating the performance of the methods MDDLSCG, MSCG and ScCG for $(m,n,k)=(256,1024,32)$
 Figure 13: Signal recovery for data $(m,n,k)=(256,1024,32)$
-Dim$(m,n,k)$ Method Itr Tcpu MSE RelErr (128,512,16) MDDLSCG 272 1.186382e-01 2.033486e-06 7.487872e-03 MSCG 493 1.606987e-01 2.033495e-06 7.488087e-03 ScCG 3653 1.135658e+00 2.033588e-06 7.712389e-03 (256,1024,32) MDDLSCG 291 2.163503e-01 2.128598e-06 9.721181e-03 MSCG 1405 8.339790e-01 2.128516e-06 9.720993e-03 ScCG 3002 1.746555e+00 2.229608e-04 9.949161e-02
+Dim(m,n,k)(m,n,k) Method Itr Tcpu MSE RelErr (128,512,16) MDDLSCG 272 1.186382e-01 2.033486e-06 7.487872e-03 MSCG 493 1.606987e-01 2.033495e-06 7.488087e-03 ScCG 3653 1.135658e+00 2.033588e-06 7.712389e-03 (256,1024,32) MDDLSCG 291 2.163503e-01 2.128598e-06 9.721181e-03 MSCG 1405 8.339790e-01 2.128516e-06 9.720993e-03 ScCG 3002 1.746555e+00 2.229608e-04 9.949161e-02
 Table 3: Comparison table for methods MDDLSCG, MSCG and ScCG
@@
 The updated method introduces spectral parameters ensuring a sufficient descent property independent of line search techniques.
-Theoretical analysis establishes the global convergence of the method for general nonlinear functions under standard assumptions.
-Numerical experiments and an application in signal processing demonstrate the improved performance and practical effectiveness of the proposed algorithm.
-Additionally, we present numerical results that validate its performance and show its superior convergence compared to the MSCG method [26] and the ScCG method [28], highlighting its potential in optimization and related applications.
+Theoretical analysis establishes the global convergence of our method for general nonlinear functions under standard assumptions.
+Numerical experiments and an application to signal processing demonstrate the improved performance and practical effectiveness of the proposed algorithm.
+Additionally, we present numerical results that validate its performance and show its superior convergence compared to the MSCG method [Faramarzi2019] and the ScCG method [Mrad2024], highlighting its potential in optimization and related applications.
 Declaration of competing interest
 The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
-Acknowledgments
-The authors are grateful to the editor and to the reviewers for their remarks and advice which improved our manuscript.
 Funding